A Complete Bibliography of The Bell System Technical Journal, 1950–1959

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 March 2020 Version 1.07

Title word cross-reference

J′_{l}(x) [Bea58]. J_{l}(x) [Bea58]. N [Gil51a, Gil58]. Z [Hel59].

-cube [Gil58]. -terminal [Gil51a].

000-Mc [LMM56].

1 [Car51, KP51].

2 [BG57, RSF51, WEP55].

3d [Mor58].

5-6 [Kin56].

60-Foot [CFJ56]. 60-kmc [Ree55].

Absorption [Bec51, Feh57]. Abstracts [Ano50a, Ano51a, Ano51b, Ano51c, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano52f, Ano53a, Ano53b, Ano53c, Ano53d, Ano53e, Ano53f, Ano54a, Ano54b]. Acceleration [Stu53]. Accounting [Kin54]. Activation [GS57]. Active [Kin59a, SDSD59]. Addendum [Rio53a]. Admittance [Edw56]. Adsorption [Kis58]. Affecting [Boz50, HE59]. algebra [Sta52]. Alloyed [EM55, ME55b]. Along [RL54]. Alphabets [Gil52, Sle56]. Alternate [Tru54]. Aluminum [BHP55, Hei51]. AM [Sun59]. Among [RFM56]. Amplifier [BM50]. Amplifiers [Ket55, LS56, MLD53, Ten56, Tho59, Tie56].

Amplitude [Rut58]. Analog [Aul59, Ble56]. Analogue [Dar51a]. Analysis [DRS56, Fuc59, HGK59, Lee55, Lou55, NA59, Pee54a, Rau59, Sum52]. Anelasticity [Kin59b]. Anniversary [Ano51j, Ano52m]. Antenna [CFJ56, Daw57, GM55b]. Apparatus [MW52]. Appearing [Ano50f]. Application [BT54, FR58, GTW53, Gra53, Hei51, Kro57, OW59, Pen58, Ung58b, Sta52]. Applications [FMW55, Hog52, HRR52, Kap55, Mil54a, Row53, SB59, Sou50, Wag54a, WP51, Dar58]. Appointed [Ano57t]. Appraisal [CR59]. approach [McC59]. Approximations [Des57]. Arbitrary [Olm58]. Arc [Ata53b, Ata53a, Ata55b, Ata55a, Ger51, SB59]. Arcing [Ata53b, Ata53a, Ata54, Ata55b, Ata55a]. Area [Sha51e]. Arising [Ric50b, Sho54]. Armature [MM51, Sum52]. Arrays [Bob57, GM55b]. Arsenide [Sha59b]. Articles [Ano50a, Ano50f]. Artificial [BHD59]. Aspects [BF59, MM57]. Assignment [Eva58]. Atlantic [GLWB57, LH57, LTLF57, MS57]. Atmospheric [CH56]. Attack [Sno57]. Attempt [Ric50b]. Attenuation [CH56, HE59, Kin56, Mar58, Mas55, Rai58, Ric57, Tho56]. Audio [TD53]. Authors [Ano50a, Ano50f, Ano50g, Ano50h, Ano51k]. Automatic [BD53, CO52, CS56, FPS56, Han56, Kin54, MV52, Pen58, Qui54, Rob56, Shi52, WEP55]. Avalanche [ME55b]. Average [Rio51]. Awarded [Ano56m].

B. [Ano57t]. B.S.T.J. [Ano59g]. Ba [Bec51]. Balanced [BP53]. Band [Loz51, Ree55, WN54]. Bands [Bul53, Daw57]. Barium [MS58]. Barrier [GS58, HCM55]. Barrier-Grid [HCM55]. Base [Lee56, TT56]. Based [Jul59]. Basic [Sch51]. Beam [DRS56, Gal59, RL54, Rig57b, Rig57a, Tie56].

Beam-Positioning [Gal59]. Beams

[RP59, Rig59]. Behavior [FMW55, KG58]. Bell [Ano50a, Ano50f, Ano50g, Ano50h, Ano51a, Ano51b, Ano51c, Ano51k, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano52f, Ano53a, Ano53b, Ano53c, Ano53d, Ano53e, Ano53f, Ano54a, Ano54b, Ano54c, Ano54d, Ano54e, Ano54f, Ano54n, Ano54o, Ano54p, Ano55a, Ano55b, Ano55c, Ano55d, Ano55e, Ano55f, Ano55p, Ano55q, Ano55r, Ano55s, Ano55t, Ano55u, Ano56a, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano57f, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano57s, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano59m, BHP55, MR50, Sha51b, Sha51e, Sha51c, Sha51d]. Bending [Fuc59]. Bends [Ung57c, Ung57b]. Bessel [Bea58]. best [SH57]. Between [Bar50, BG55, Tie56]. Beyond [CHK59, FCH57]. Binary [DeL56, GM59, Llo57, Row58, Sle56, Sun59, Ulr57, Wra56, Hag59, Lee59]. binary-decision [Lee59]. binomial [SH57, SG59]. Biological [Sno57]. Bivariate [OW59]. Block [Col53a, Llo57]. Blocking [Clo53, NA59]. Bombarded [KO52, Ohl52]. Boolean [HS55, McC56a, McC56b, Sta52]. Borocarbon [GPV51]. Break [Ata54]. Breakdown [FR58]. Breaking [WM59]. Bridges [Wil52]. Brillouin [RP59]. Broadband [BDS58, Jac58]. Buckled [Hop52]. Buildings [Ric59]. Bulloch [Ano57t]. burst [Hag59]. burst-correcting [Hag59].

C [Kel51]. C. [Ano51i, Dar51b]. Cable [BHP55, EH57, Epp54, Gil51b, Hop52, JLL57, LFB57, MMVH57, MHEG57, Sno57, Win53, Woj54, Zaj57]. Cables [Mau51, Rai58, RL55, Sha51e]. Calling [MPW58]. Calls [Rio53b, Wil53, Rio53a]. Campbell [Ano55o]. Capacitance [CS56, Lee58, McC57]. Capacitors [CS56]. Carbon [GPV51]. Card [BT54, HN53]. Carrier [BHP57, Car51, Ear55, EM52, HRR52, KP51, Lev58]. Cathode [Jen51, TD53, Tow57]. Cavities [SB58]. Central [Ell59, YCU58]. Centralized [Kin54]. Certain [McC57]. Chain [Row58]. Channel [Bab53, PK57, WEP55, Bur59, Sha59a]. Channels [Loz51, MM57]. Character [Sei57]. Characteristic [Kle56]. Characteristics [Ata53a, Bar50, FM56, MS58, Pee54a, RM58, War58]. Charge [RP59, Rig59, Tie56]. Chart [Rob58]. Chemical [RFM56]. Circuit [Ben53a, KG58, Tie56, WR51, WP51, YCU58, Sta52]. Circuits [And52, Ata53b, Ata53a, Ata54, Ata55a, Ble56, Han56, HV55, Mea55, Pie51b, Gil51a, HS55, Lee59]. Circular [Kin56, Ung57a, Ung57b, Ung57c, Ung58a, War58, Mor57]. Class [Sle56]. Clogston [Kin59c, Vaa53]. Closing [Ell55]. coated [Ung57a]. Coaxial [Ano53m, DKK53, Ear55, EEKG53, FW51, GTW53, Kem51, KF53, MLD53, Rai58, RM58, RG53]. Code [Goo51]. Codes [Gil58, Ham50, Ulr57, Hag59, McC59, Sha59a]. Coding [Jul59, Llo57, Oli52]. Coefficient [Mas51]. Coefficients [McC57]. Coincidences [GP57]. Cold [Hei51, TD53, Tow57]. combinational [HS55]. Combined [BG56]. Combining [Har59]. Common [Mye52]. Communication [BHW59, Ric50a, Vau59]. Communications [BT58a, BT58b, KR57, Mil54b]. Companding [Smi57]. Compandors [Loz51, Mal51]. Comparators [NM59]. Comparison [Gil52, You52]. Completely [SW54a]. Components [GTW53]. computer [LLM59]. Computing [OW59]. Concentrated [Egg50]. Concentrating [Joh54]. Concentration [Bac58, Bar50].

Concentrator [Joe56b]. Concepts [HSK58, Sch51]. concerning [Pri51]. Conditions [BCR55, KG58, MO53]. Conductive [Woj54]. Conductivity [VP57]. Conductivity-Modulated [VP57]. Conductor [GM53, RM58, Uhl55]. Conductors [Clo51, Vaa53]. Conduit [Wei57, WM59]. Configuration [Pri53]. Connection [MO53, MA54, Pri57]. Connections [Ell59, Mal53, MO53, McR53, Van53]. Considerations [Feh57, Tho59]. Constant [Dar50, Mau51, Bur59]. Constants [BMM51]. Contact [AC54, BP53, FM56, KG58, Kis58, Qui54, Ric50b, Sha59b, SB59]. Contacts [Ata53b, Ata53a, Ata54, AC54, Ata55b, Ata55a, Bar50, Ell55, GS57, HE58]. Containing [Des57, Mor57]. Continuous [Mur58, Woj54, Ben53b]. Continuously [Rai58]. Contribution [Lev58]. Contributors [Ano50b, Ano50c, Ano50d, Ano50e, Ano51d, Ano51e, Ano51f, Ano51g, Ano51h, Ano52g, Ano52h, Ano52i, Ano52j, Ano52k, Ano52l, Ano53g, Ano53h, Ano53i, Ano53j, Ano53k, Ano53l, Ano54g, Ano54h, Ano54i, Ano54j, Ano54k, Ano54l, Ano55g, Ano55h, Ano55i, Ano55j, Ano55k, Ano55l, Ano56g, Ano56h, Ano56i, Ano56j, Ano56k, Ano56l, Ano57g, Ano57h, Ano57i, Ano57j, Ano57k, Ano57l, Ano58a, Ano58b, Ano58c, Ano58d, Ano58e, Ano58f, Ano59a, Ano59b, Ano59c, Ano59d, Ano59e, Ano59f]. Control [Bab53, DKK53, Epp54, GTW53, Log54, Mye52, Pee54b, Rob58, Sch57, YCU58]. Controlled [Joe56b, MV52, Rau59]. Controls [BH55]. Conversion [Aul59, Edw56, Egg50, Lec59, Sch55]. Converter [Lee58, Mer51]. Converters [Uhl58]. Copper [Lec59]. Core [Sto53]. Correcting [Ham50, Hag59, McC59]. Correction [Ano55m, Ano55n, Ulr57]. Correlation [Col53a, Ben53b]. correlatograph [Ben53b]. Cost [Eva58]. Counts [Hay52]. Coupled [CKQ56, KM56, Mil54a, Ree53]. Coupler [Lou55]. Couplers [Coo55]. Coupling [Fox55, Tie56]. Cr [GPSD59]. Criterion [Mur58]. Cross [GB56]. Cross-Sections [GB56]. Crossarms [Egg50]. Crossbar [Ada56, BD53, ME55a, Pen58]. Crosstalk [BRT56]. Crystal [MS58]. Crystals [BS56a, Mac51, Mas50, Mas51, Woo51]. cube [Gil58]. Current [Blo53]. Curved [Mar58, Mor57]. Curves [Rio53b, Wil53, Rio53a]. Customers [CW52]. Cut [Olm58]. Cylinder [Ric54]. cylinders [Hol51]. Cylindrical [RL54, RP59, SW54a].

D [Ano57t]. Data [MM57]. Davisson [Ano51i, Dar51b, Jen51, Kel51]. DC [Pri53]. decision [Lee59]. defectives [SG59]. Defects [RFM56]. Definition [SH58]. Deflection [Aul59, Jen51]. Delay [Rio53b, Bur59, Rio53a]. Delayed [Wil53]. Dependence [BMM51]. Deposits [HE58]. describing [Pri51]. Design [Ano54m, BWW57, Ear53, EM55, EEKG53, Fry52, Fuc59, GM55b, GLWB57, HB57, HSK58, HHH59, Hun53, KP51, Kro57, LFB57, LTLF57, LS56, Log54, MR50, NA59, PW54, Pee54c, Pfe54, SY58, Tho59, Ung58a, WR51, YCU58, HS55, Sta52]. Designed [BT54]. Detailed [DRS56]. Detecting [Ham50]. Detection [Jul59, Kap55, McC56a]. Deterioration [Big51]. Determination [Fuc59, McC57]. Determined [Olm58]. Determining [Tru54]. Detwinning [Woo51]. Development [FW51, GM53, HP55, Lev58, Mor52c]. Developments [GGI58]. Device [GC59]. Devices [FR58, Gru59, Log53]. Dialing [CW52, FL56, HN53]. Dials [Pfe54]. Diameter [CFJ56]. Dielectric [Mau51, MS58, VR57, Mor57, Ung57a]. dielectric-coated [Ung57a]. Difference [Dar50]. differential [Pri51]. Diffraction [Ric54]. Diffused [Bac58, Lee56, Pri56, TT56]. Diffusional [Ric50b]. Digital [Aul59, Ben58, Ble56, HCM55, HV55, Ket59, NM59, LLM59]. Digital-to-Analog [Aul59]. Diode [Gra51, Hus53, Kle56, Rea58]. Diodes [CHS54, Sho54]. Direct [FL56]. Directive [GM55b]. Discharge [Ata54]. Discs [SB58]. Dislocations [Mas55]. Displacement [WS56]. display [Ben53b]. Displaying [KH51]. Dissipated [Ell55]. Dissipation [Ger51]. Distance [Ada56, FL56]. Distortion [Ket55, Ric57]. Distributed [Egg50]. Distribution [GB56, Pri53, Ric58, Bur59]. Distributor [Fre58]. diverging [Har50b]. Domain [Gal54, Gal55]. Dr. [Kel51]. Drag [HGK59]. Drop [GM53]. Drum [BHM56]. Dual [Daw57, ME55a]. Duality [WR51]. Due [CR57, KM56, PW56]. duration [Ric58]. Dynamic [Log53]. Dynamics [Zaj57].

Each [Egg50]. easily [Hag59]. Economics [Wag54a]. Economy [Bul53]. Edge [Jul59]. Editor [Ano57t, Ano59g]. Editorial [Bak58]. Effect [BG56, Clo51, DR55, FM56, Gru59, Mas55, Hog52]. Effects [HGK59, Mas50, Stu53, Suh53, Tie56, War58, WM59]. Efficient [Oli52]. efficiently [SG59]. Effort [Mur58]. Elastic [BMM51, Hol51]. Electric [Ger51, Kin56, Ung57c, Wan53, War58, Ung57a, Ung57b]. Electrical [Ata53b, Ata53a, Ata54, Ata55b, Ata55a, GS57, KHD55, Mon52]. Electrically [Sch57]. Electro [Mas50]. Electro-optic [Mas50]. Electrodes [Ger51]. Electroformed [GM53]. Electrolyte [BG55]. Electrolytic [Uhl56]. Electromagnetic [Log53]. Electron [Bec51, Cal51, DRS56, FW51, Gra51, Hei51, MMVH57, Pie51b, RL54, Rig57b, Rig57a, Rig59, Stu53]. Electronic [BH55, DB54, FG58, Fre58, Joe58, SY58, YCU58]. Electronically [MV52]. Electronics [Joe56a]. Electrons [Dar53b, Mac51, Pie51a, Sho51, Van50, Pri51]. Element [Ban54, Bob57]. Elements [GM55a]. eliminate [SG59]. Emerald [GPSD59]. Emission [Bec51]. Emitter [TT56]. Encoding [Ric50a]. Encodings [GM59]. Energy [Ell55]. Engineering [BT58a, BT58b, BF59, EH57, HE59, Tru54, Uhl55, Wil56]. English [Sha51a]. Entropy [Sha51a]. Equalization [KF53, RL55]. Equalizer [BG57]. Equations [Sch52, Pri51, Sch55]. Equilibrium [Bur59]. Equipment [BD53, Joh54, MS57, Wag54b]. Eroding [KG58]. Erosion [Ata55a, SB59]. Errata [Ano57m]. Error [Ham50, McC59, Ric50a, Ulr57, Sha59a]. Error-correcting [McC59]. Errors [OW59]. Estimation [Bur58, Log54, Pee54b]. Evaluation [Bac58, Col53a, Col53b, Ell59, Van53]. Evolution [Mar51, Sha51b, Sha51e, Sha51c, Sha51d]. Exceptional [BS56a]. Exchange [Ben57b, Sha50, Sha51e]. Excitation [RP59]. Experience [Col53a]. experiment [Sob56]. Experimental [BHD59, BHW59, BG57, BHP55, Daw57, Joe56b, Joe58, Kin59c, LMM55, MV52, LLM59]. Experiments [BG55, BT54, DeL56, DP58, Har52]. Exponential [Wil53, Ben59]. Expressed [Mas50]. Extremely [BS56b]. Extrusion [BHP55].

Facilities [Sha51b, Sha51e, Sha51c, Sha51d]. Factors [Boz50, Fry52, Gra53, MR50]. fades [Ric58]. Fading [CJ52, Kay53]. Faraday [Hog52]. Fast [GB56]. Features [Car51, KP51]. Feed [Daw57, MS57, TK57]. Feedback [Ket55]. Ferrite [Gal54, Gal55, SB58, SF59, Sto53]. Ferrites [Clo55, FMW55, Row53]. Ferroelectric [Woo51]. Ferromagnetic [Hog52, VR57]. Field [Bec51, BG56, DR55, Pie51a, Pri53, Rig57b, Rig57a, WS56]. Fields [Wan53]. Fifth [Ano51j]. Figure [Lee58]. Filament [Suh53]. Filled [SW54a]. Film [GPV51]. Filters [War58]. Finite [Tie56]. First [Bak58, Bea58]. Five [Mar51]. Flow [Van50, Pri51]. Fluctuation [Ben57a]. Fluctuations [Ric50b, Sle58]. Flying [Gal59, HSK58, HHH59, PDH59]. FM [BCR55, Ric57, Rut58, Sun59]. Focused [RL54]. Foot [CFJ56]. Forces [MW52]. Foreword [Ano53m]. Form [Mas50]. Formal [McM52a, McM52b]. Formation [DRS56]. Formulae [Fuc59]. Forward [Kle56]. Four [Smi58]. Four-Point [Smi58]. Fourth [Pie50b]. Frequencies [De 52, Hog52, MM52, VR57]. Frequency [Bab53, BMM51, Bul53, Edw56, Kay53, Lee56, Rea58, SB58, Sha59b, Tho59, TD53, Uhl58, WN54]. function [McC56a]. Functional [SY58]. Functions [Bea58, FPS56, McC56b, Sem58, Zin52]. Fundamental [HSK58, Pil52, SB59, Sum54]. Fundamentals [Bul57, Sun54a, Sun54b].

Gain [Lee58]. Gallium [Sha59b]. Gas [Tow57]. Gaseous [Wan53]. Gases [Kis58]. Gates [Hus53]. Gaussian [Ric58, Sha59a]. General [Kel52]. Generalizations [Pri57]. Generalized [Sch52, Sch55]. Geometries [McC57]. George [Ano55o]. Germanium [Bar50, BB53, BG55, EM55, GB56, HGK59, Lee56, RFM56, Sho51, Suh53, Uhl56, Van50]. Governor [Pfe54]. Gray [Gil58]. Grid [GS58, HCM55]. Group [SG59, McC56a]. Growing [PW56, Rig57b]. Grown [ATS59]. Guest [Bak58]. Guide [SW54a, WR51]. Guided [SW54a, SW54b, SW54c]. Guided-Wave [SW54a]. Gun [Pie51a]. Guns [DRS56]. Gyrator [Hog52]. Gyromagnetic [Sei57, SF59, SW54a, SW54b, SW54c]. gyrostatic [Har50b].

H. [Ano59g]. Habits [CW52]. Hall [Gru59].

Handling [OW59]. Harmonics [Rig59]. Haul [HRR52]. Heat [Ger51]. Helices [CKQ56]. Helix [Lec59, Mar58, MY56, Ros58, SW54b, Ung58b]. High [Aul59, De 58, EM55, FG58, GS58, Kay53, Lee56, Mas51, NM59, Rea58, Sha59b, Tho59, VP57]. High-Frequency [Lee56, Rea58, Sha59b, Tho59]. High-Speed [Aul59, De 58, EM55, FG58, GS58, NM59]. High-Voltage [VP57]. Historical [KR57]. holding [Ben59, Bur59]. Hole [Bar50]. Holes [Van50, Pri51]. Horizon [CHK59, FCH57]. Hot [Sho51]. Human [PK57]. Hydraulic [Rau59, Sch57].

Ideal [Sun59]. Identification [Pen58]. Identifiers [Sch51]. II [Ata53a, BT58b, Col53b, Dar53b, Log54, MO53, McM52b, Mor52b, Rig57a, SW54b, Sun54b]. III [Ata54, Sha51e, SW54c, Van53]. Impedance [GM55a, Kem51, Mer51, MRS54, Wil52, Zin52]. Implications [HE59]. Important [Fry52]. Improved [Ben53a, IT51, Ung58a]. In-Band [WN54]. Including [NA59, Tie56]. Incremental [Woj54]. Individual [Gal54, Gal55]. Inductive [Ata54, Sha51b, Sha51e, Sha51c, Sha51d]. Inductors [Sto53]. Infinite [Tho56, Uhl55]. Influencing [Fry52]. Information [EH57, HV55, Kel56, PK57]. Inherent [NA59]. inhomogeneous [Mor57]. Initiation [Ata53b, Ata55b]. Inorganic [Cal51]. input [Bur59]. Inspection [Mur58]. Installment [Pie50c, Pie50d, Pie50b]. Instantaneous [Loz51, Mal51, Smi57]. Insulated [Win53]. Insulation [BD53, KHD55]. Integrated [Vau59]. Interaction [BH52]. Interactions [RFM56]. Interaxial [Mau51]. Interchannel [BCR55, CR57]. Interface [BG55]. Interference [Bab53, BCR55, BS56b, CR57]. Intermodulation [Bab53]. Internal [Pee51]. Interpretation [Col53a, Kel56].

Intertoll [FPS56, FL56, Hun53, Joh54]. Intrinsic [Pri53, Rea56]. Introduction [Ano54m, McM52a, McM52b, McR53]. invariance [McC56a]. Inventors [Ano56m]. Investigations [Bec56]. Ionic [Ohl52]. Ionically [KO52]. Ions [Wan53]. Iron [Gal54, Gal55]. Isolator [WS56]. isotropic [Hol51]. Issue [Ano50b, Ano50c, Ano50d, Ano50e, Ano51d, Ano51e, Ano51f, Ano51g, Ano51h, Ano52g, Ano52h, Ano52i, Ano52j, Ano52k, Ano52l, Ano53g, Ano53h, Ano53i, Ano53j, Ano53k, Ano53l, Ano54g, Ano54h, Ano54i, Ano54j, Ano54k, Ano54l, Ano55g, Ano55h, Ano55i, Ano55j, Ano55k, Ano55l, Ano56g, Ano56h, Ano56i, Ano56j, Ano56k, Ano56l, Ano57g, Ano57h, Ano57i, Ano57j, Ano57k, Ano57l, Ano58a, Ano58b, Ano58c, Ano58d, Ano58e, Ano58f, Ano59a, Ano59b, Ano59c, Ano59d, Ano59e, Ano59f]. IV [Ata55b].

J [Ano51i, Dar51b, Kel51]. J. [Ano57t]. Journal [Ano51a, Ano51b, Ano51c, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano52f, Ano53a, Ano53b, Ano53c, Ano53d, Ano53e, Ano53f, Ano54a, Ano54b, Ano54c, Ano54d, Ano54e, Ano54f, Ano54n, Ano54o, Ano54p, Ano55a, Ano55b, Ano55c, Ano55d, Ano55e, Ano55f, Ano55p, Ano55q, Ano55r, Ano55s, Ano55t, Ano55u, Ano56a, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano57f, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano57s, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano59m, Ano50f, Ano50g, Ano50h, Ano51k]. Junction [CHS54, Ear53, Ear54, EM55, Lec59, ME55b, Pri56, Uhl58].

Kinematics [Zaj57]. Klystron [CR57, Ree53, Ree55]. kmc [Har59, HE59, Ree55].

L3 [Ano53m, DKK53, Ear55, EEKG53, GTW53, KF53, Lev58, MLD53, RG53]. Laboratories [BHP55, MR50]. Laboratory [BHM56, Col53a]. Laddic [GC59]. Laminated [Clo51, Mor52b, Mor52a, Rai59, Vaa53]. Large [Bob57, FR58, Tie56]. Large-Sized [Bob57]. Larynx [BHD59]. Law [Sho51]. Layers [Bac58, BS56b]. Laying [JLL57, Zaj57]. least [Dar58]. least-squares [Dar58]. Length [GM59]. Level [DSDS59, SDSD59]. Leveling [BS56a]. Limit [Mur58]. Line [BD53, FG58, Jac58, Joe56b, Kin59c, MM57, RM58]. Linear [Dar58, Har52, Ric50b, McC59]. Lines [BRT56, GM55a, Mor52b, Mor52a, Rai59]. Link [BWW57, HB57, TK57, GLWB57, LH57, LTLF57, MS57]. Load [Hay52]. Loaded [Rai58, SF59]. Loading [BCR55, Sha51b, Sha51e, Sha51c, Sha51d]. Loads [Egg50]. Local [BD53, DK55, RL55]. Logic [GC59, Ket59, NM59, LLM59]. Long [Ada56, Row58]. Longitudal [Rig57a]. Longitudinal [Hol51, Rig57b]. Longitudinally [SB58]. Loosely [KM56]. Loosely-Coupled [KM56]. Loss [FL56, KM56]. Losses [BMM51, Clo51]. Low [Kin59b, Ree55]. Low-Voltage [Ree55].

Machine [CS56, FKR53, Ben53b]. Magnetic [All53, Boz50, BHM56, BHN55, Dar53a, Dar53b, GC59, PW54, Pie51a, Rig57b, Rig57a, Suh53, VR57]. Magnetically [RL54, Wal51]. Magnetically-Focused [RL54]. Magnetization [Pee54a, SW54b]. Magnetized [SB58]. Maintenance [FL56]. Manufacture [Epp54, GTW53, GM53, LFB57, Qui54, Rob56]. Manufacturing [Ear55, Eva58, Han56]. Many [Her55]. Many-Valley [Her55]. Marine [Sno57]. Maser [DSDS59]. Masers [SDSD59]. Material [SDSD59]. Materials [Col53a, Sno57, VR57]. Mathematical [Mor52b, Mor52a]. matrices [HS55]. Matter [Har50a]. Maxwell [Sch55]. mc [Ric59, LMM56, You52]. Means [De 52, Edw56]. Measure [Kis58]. Measured [Pee54a]. Measurement [BT58a, BT58b, BS56b, CH56, Ell55, Epp54, Smi58, VR57]. Measurements [BG56, Hay52, Log53, Woj54]. Measuring [MW52, OW59, Wag54b]. Mechanical [BH52, MM52]. Mechanism [Ata55b]. Mechanisms [Ata55a, Ric50b]. mechanized [Hag59]. Media [Sei57, SW54a, SW54b, SW54c]. Medium [Kin56, LMM56, Men52, Mil54b, Har50b]. Medium-Sized [Kin56]. Megacycle [Wil52]. Memory [Bob57, HCM55, Sha50]. Mercury [BP53]. Message [Kin54, Men52]. Metal [Hei51, HE58]. Metallic [Mac51]. Metals [Kis58, Mas55, Mor58]. Method [Dar51a, Jul59, KH51, Mea55, OW59]. Microscope [Bec51, BS56b]. Microscopy [Cal51]. Microwave [BM50, De 52, DeL56, Feh57, FMW55, Hin50, Hog52, KH51, Pie54, RSF51, Row53, VR57]. Microwaves [CJ52]. Migration [KHD55]. Military [GGI58]. Millimeter [CH56, Sha56]. Millimicrosecond [Bec56]. Minimization [McC56b]. Minimum [Eva58]. Miscellaneous [SW54c]. mm [Kin56]. Mobile [Bul53, You52]. Mode [Har50a, KM56, Lec59, Lou55, Ung57c, War58]. Model [Ben57b, BHM56, Vau59, Ric58]. Models [Ros58]. Modes [Fox55, KM56, Sei57, SF59, Hol51]. Modulated [Ric57, VP57]. Modulation [Goo51, Jen51, Rut58]. Modulator [BM50]. Modulators [BHN55]. Molding [BCS54]. Monographs [Ano54n, Ano54o, Ano54p, Ano55p, Ano55q, Ano55r, Ano55s, Ano55t, Ano55u, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano57s, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano59m]. Motion [Gal54, Gal55, Har50a, Wan53]. Multi [KM56]. Multi-Mode [KM56]. Multicontact [Raf54]. Multifrequency [Blo53]. Multipaired [Mau51]. Multistage [FR58].

N [Ear54, HGK59, Pri56, Uhl58, WP51, Car51, Ear54, KP51]. N-1 [Car51, KP51]. N-P-I-N [Ear54]. n-Type [HGK59]. Narrow [Loz51]. Nationwide [CO52, HN53, Nun52]. Natural [Kin59b]. Nature [Cut56, She58]. Necessary [MO53]. Negative [GM55a, Mer51, MRS54, Rea58, Sho54, Ben59]. Negative-Resistance [Rea58]. Net [FL56]. Network [Ban54, BH55, Dar51a, Dar52, Des57, Har59, Sem58, Zin52]. Networks [Clo53, CS56, FR58, Kin59a, Lee55, Pri57, Tru54]. Newfoundland [BWW57, HB57, TK57]. Nickel [BMM51, Gal54, Gal55]. No [Pen58]. Nobel [Ano56m]. Noise [BCR55, Lee58, Mon52, Ric50a, Ric57, Ric50b, Rig57b, Rig57a, Rig59, Sle58, Suh53, Uhl58, Ric58]. Non [Clo53, KG58, SW54b, Ulr57, Woj54]. Non-Binary [Ulr57]. Non-Blocking [Clo53]. Non-Conductive [Woj54]. Non-Eroding [KG58]. Non-Reciprocal [SW54b]. Nonlinear [Edw56, Ric57]. Nonlinearities [NA59]. Nonparametric [SH58]. Nonstationary [Bur58]. Nonuniform [Jac58]. Nonuniformities [Rai59]. Normal [Fox55, OW59, Ung57c]. North [GLWB57, LH57, LTLF57, MS57]. Nova [BWW57, HB57, TK57]. Nuclear [Dar53a]. Number [Pen58]. Numbering [Nun52]. numerical [Uhl55].

O [EM52]. Objectives [Car51]. Observed [Kin56]. Obtaining [MO53]. Occurrence [Bab53]. Oceanographic [EH57]. Office [Ell59]. Offices [BHM56, Pen58]. Ohm [Sho51]. Oil [Col53a]. Oil-Type [Col53a]. one [Bur59, SH57]. Open [AC54, BRT56].

Open-Contact [AC54]. Open-Wire [BRT56]. Operate [Log54, Pee54b]. Operated [Des57, Sch57]. Operation [ME55a, Ree55]. Operational [FPS56]. Operator [FL56]. optic [Mas50]. Optical [Mas50]. Optics [PDH59]. optimal [Sha59a]. Optimum [Eva58, GM55b, Log54]. Order [Wil53]. Organic [Big51, GS57, HE58, Sno57]. Origin [She58]. Oscillator [Blo53, Hin50, NA59]. Oscilloscope [BDS58]. Other [Ano50g, Ano50h, Ano51k, Hop52, Van50]. Outdoor [Col53a]. Over-All [MHEG57]. Oxides [ATS59, Mor58].

P [Ear54, Pri56, Uhl58, WP51, Ear54]. P-N-I-P [Ear54]. Pairs [Mau51]. Panel [Pen58]. Papers [Ano51a, Ano51b, Ano51c, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano52f, Ano53a, Ano53b, Ano53c, Ano53d, Ano53e, Ano53f, Ano54a, Ano54b, Ano54c, Ano54d, Ano54e, Ano54f, Ano54n, Ano54o, Ano54p, Ano55a, Ano55b, Ano55c, Ano55d, Ano55e, Ano55f, Ano55p, Ano55q, Ano55r, Ano55s, Ano55t, Ano55u, Ano56a, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano57f, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano57s, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano59m]. Parabolic [CFJ56, Ric54]. Parallel [Sem58]. Paramagnetic [Feh57, GPSD59, SD59]. Part [Ata53b, Ata53a, Ata54, Ata55b, Ata55a, BT58a, BT58b, Col53a, Col53b, Dar53a, Dar53b, Log54, Mal53, MO53, Mor52b, Mor52a, Pee54b, Rig57b, Rig57a, SD59, Sha51e, SW54a, SW54b, SW54c, Van53]. partial [Pri51]. Partially [SF59]. Paths [Gil58]. Patterns [GP57, KH51]. Pavement [WM59]. Percussion [SB59]. Percussive [Sum54]. Perfection [BS56a].

Performance [AC54, BM50, Kro57, MHEG57, Van53]. Performing [GC59]. Periodically [Des57]. Permanence [MA54]. Permanent [MO53]. Perturbation [SW54c]. Phase [Dar50, Ric57, Tho56]. Phenomena [All53, Ata54, Clo55]. Phenomenon [Rig57b]. Philosophy [YCU58]. Phonon [HGK59]. Phonon-Drag [HGK59]. Photo [BG56]. Photo-Voltage [BG56]. Photoconductivity [BG56]. Photoelastic [Mas50]. Photoelectric [KO52]. Photographic [KH51]. Photography [PDH59]. Physics [Ano56m]. Picture [Pie54]. Pierce [DRS56]. Piezoelectric [MS58]. Pin [Egg50, Kle56]. Pipe [Lec59]. Pipes [Kin56]. Plan [BCH55, Nun52]. Plane [Ric54]. Planning [MHEG57]. Plans [Mur58, Pil52]. Plant [Pil52]. Plants [CHS54]. Plates [MS58]. Plot [Col53a]. PM [BCR55]. Point [Bar50, BT58a, BT58b, FM56, Sha59b, Smi58]. Point-Contact [FM56, Sha59b]. Poisson [Bur59, GP57]. Polar [BP53]. Polarization [Daw57]. Polyethylene [Epp54, Win53]. Polymers [BH52, Big51, MM52]. polynomial [Dar52]. Polytonic [LMM55]. populations [SH57]. Position [Egg50]. Positioning [Gal59]. Potential [Dar51a]. potentials [Uhl55]. Power [BT58a, BT58b, CHS54, Cut56, Ell55, LMM56, MS57, Sle58, TK57]. Power-Feed [TK57]. Precious [HE58]. Precision [Ear55]. Prediction [Har52, Sha51a, Dar58]. Presence [Ric50a]. Present [Mor52c]. Preservatives [Col53a, Col53b]. Pressure [McC57]. Principle [Hop52]. Principles [Pee54c, Pfe54, Sou50]. Printed [Sha51a]. Private [MM57]. Prize [Ano56m]. Probabilities [OW59]. Probability [Ric50a, Sha59a]. Probe [Smi58]. Problems [Sum54, Uhl55]. process [Ben59]. Processes [SB59]. Produced [Ric57]. Production [LH57]. Program [Lev58, SY58]. programming [McC59]. programs [Lee59]. Propagation [Bul57, CFJ56, CHK59, De 52, RL54, Rig59, SW54a, SW54b, SW54c, FCH57]. Properties [BB53, Her55, KO52, Mas50, MM52, Ohl52, Rob58, Vaa53, VR57, WP51]. Proposed [Rea58]. Protection [WEP55]. Publications [Ano50g, Ano50h, Ano51k]. Published [Ano51a, Ano51b, Ano51c, Ano51i, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano52f, Ano53a, Ano53b, Ano53c, Ano53d, Ano53e, Ano53f, Ano54a, Ano54b, Ano54c, Ano54d, Ano54e, Ano54f, Ano54n, Ano54o, Ano54p, Ano55a, Ano55b, Ano55c, Ano55d, Ano55e, Ano55f, Ano55p, Ano55q, Ano55r, Ano55s, Ano55t, Ano55u, Ano56a, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano57f, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano57s, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano59m]. Pull [Pee54a]. Pulling [CR57]. Pulse [BHN55, Goo51, Sun54a, Sun54b, Sun59, Ten56, Wra56]. Pulses [Bec56, De 52, DeL56]. Pulsing [Blo53]. Purpose [Kel52]. Pushbutton [MPW58]. Pyrolytic [GPV51].

Quality [Boz50, DKK53, GTW53, She58]. Quantized [Smi57]. Quartz [Kin59b, Mas51].

Radio [Bab53, BG57, Bul53, Bul57, Daw57, Har59, HE59, Kay53, LMM56, Ric54, Ric59, RSF51, WEP55, You52, Ric58]. Random [GM55b, Rio53b, Sle58, Wil53, Bur59, Rio53a]. Range [Hin50, Wil52]. Rate [Kel56, PK57]. Rates [PK57]. Rating [BCH55]. Ray [Jen51]. RC [Kin59a]. Reading [PK57]. Realizability [McM52a, McM52b]. Realization [Dar50]. Rebound [Sum52]. Received [CR59].

Receiver [MM51]. Reciprocal [SW54b]. Recorded [Wal51]. Recording [Men52]. Recovery [Hei51, Zaj57]. Rectifier [VP57]. Rectifiers [Gra53, Pri56, Sha56, Sha59b]. Recurrent [Hag59]. reducing [Sob56]. Reduction [Clo51]. Reed [HP55]. Reference [Col53a]. Reflection [Mac51, FCH57, Har50b]. Reflex [Ree53, Ree55]. Regeneration [DeL56]. Regenerative [Ben58, DP58, De 58, Row58, Sun57, Ten56]. Regenerator [Wra56]. Region [FMW55]. Regulation [KF53]. Relation [Bar50]. Relaxation [Clo55]. Relay [BG57, BP53, DB54, Daw57, Han56, HE59, KG58, Kel52, LMM56, Pee51, Pee54c, Qui54, Raf54, RSF51, Sum52, Wag54a, Wag54b, WEP55, HS55]. Relays [Ano54m, BCS54, Fry52, HP55, Log54, ME55a, Pee54b, PW54]. Release [Pee54c]. Reliability [Fry52, Hay52, Sob56]. Remote [Joe56b]. renewal [Ben59]. Renne [Ano59g]. Repeater [BWW57, GLWB57, LH57]. Repeaters [DP58, De 58, Gil51b, MRS54, Row58, Sun57]. Replication [Cal51]. Representation [Lee59, Zin52]. Representativeness [SH58]. Reproduction [Wal51]. Requirements [DKK53, GTW53, Sha50, Tru54, Wei57]. Research [Ros58, Vau59]. Resistance [CS56, Kis58, Rea58, Sho54, Sno57]. Resistance-Capacitance [CS56]. Resistivities [Smi58]. Resistors [GPV51]. Resolver [Kro57]. Resonance [Dar53a, Dar53b, Feh57, GPSD59, KM56]. Resonator [Ree53]. Results [Col53a, SW54c, Pri51]. Resume [KR57]. Review [All53]. Revised [BCH55]. Ring [MM51]. Ringing [MPW58]. Rotating [Rig59]. Round [Wei57, WM59]. Route [JLL57]. Routing [Tru54]. Ruby [SD59]. Runs [Olm58]. Rural [BHP57].

S [Ano57t, Ano59g]. Sample [Olm58, SH58, SG59]. Sampling [Mur58]. Sapphires [SD59]. Saturation [Cut56]. Scanner [FG58]. Schemes [Ric50a]. Scientific [Dar51b]. Scotia [BWW57, HB57, TK57]. Second [Pie50c]. Sections [GB56, Hei51]. Selecting [SH57]. Selection [Bab53, JLL57]. Selective [CJ52, Kay53]. Selenium [Gra53]. Self [Hei51, Sun57]. Self-Recovery [Hei51]. Self-Timing [Sun57]. Semi [Tho56, Uhl55]. semi-conductor [Uhl55]. Semi-Infinite [Tho56]. Semiconductor [Her55, Hus53, Pri53, Sho54, YCU58]. Semiconductors [Mon52, Van50, Pri51]. Sensitivity [Feh57]. Separation [Vau59]. Sequential [Mea55]. Series [Sem58, Dar52]. Series-Parallel [Sem58]. serpentine [Ung57b]. Served [Rio53b, Wil53, Rio53a]. Service [BHP57, CO52, MM57, WEP55, Bur59]. Services [Men52]. serving [Ben59]. Servo [Gal59, Ket59]. Servomechanisms [Rau59]. Set [Ben57b, Ben53a, IT51]. Seventy [Ano51j, Mar51]. Seventy-Fifth [Ano51j]. Seventy-Five [Mar51]. several [SH57]. Shaping [Uhl56]. Sheath [BHP55, Epp54, Woj54]. Sheaths [Hop52]. Sheet [Smi58]. Shift [Ric57]. Shifts [SB58]. Short [Ata53b, Ata53a, Ata55b, Ata55a, De 52, GM55a, Ger51, HRR52, SB59, Ben53b]. Short-Haul [HRR52]. Shortest [Pri57]. Shot [Uhl58]. Signal [Fre58, Kap55, Ric57, Tie56]. Signal-Detection [Kap55]. Signaling [LMM55, Sle56, WN54]. Signalling [Gil52]. Signals [Jul59, Kay53, Kre52, Smi57, TD53, Wal51]. Silicon [ATS59, Bac58, KO52, Ohl52, Pri56, RFM56, TT56, Uhl56, VP57]. Silver [KHD55]. Simple [Ben57b]. Single [BS56a, Lou55, MS58, WN54]. Single-Crystal [MS58]. Single-Frequency [WN54]. Sized [Bob57, Kin56]. Skin [Clo51]. slabs [Hol51]. Slope [Tho56]. Slow [Pee54c]. Small [Kin56, SB58]. smoothing [Dar58]. Soil [Col53a]. Soil-Block [Col53a]. Solderless [Ell59, Mal53, MO53, MA54, McR53, Van53]. Solid [DSDS59, RP59, SDSD59]. Solid-Cylindrical [RP59]. solutions [Uhl55]. Solved [Des57]. Some [HE59, KP51, NM59, Pri51, Pri57, Sch51, Sum54, Tho59, WP51]. Sound [KH51]. sources [Uhl55]. Space [KH51, RP59, Rig59, Tie56]. Space-Charge [RP59, Rig59]. Spacing [Mau51]. Special [Col53a]. Spectra [BT58a, BT58b, SD59]. Spectrum [GPSD59, Rig57b, Rig57a]. Speech [CR59, Loz51]. Speed [Aul59, De 58, EM55, FG58, GS58, NM59]. Spot [Gal59, HSK58, HHH59, PDH59]. Spring [BCS54, Qui54]. squares [Dar58]. Stability [GM55a]. Stabilization [ATS59]. Stable [Aul59]. Standards [Kem51, She58]. Starting [Pie51a]. State [DSDS59, SDSD59]. States [GB56]. Statistical [Kay53, Sob56]. Statistics [Ben57b, Ben58, Kre52, Lev58]. Status [Mor52c]. Step [Pen58]. Step-by [Pen58]. Storage [Aul59, Bob57, HCM55]. Store [Gal59, GS58, HSK58, HHH59, PDH59]. Stored [SY58]. Stored-Program [SY58]. Straight [Pie51a]. Straightening [BCS54]. Strains [Hop52]. Streams [Gra51, Pie51b]. Strength [Wei57]. Stress [MA54]. Strong [Wan53]. Structure [Mal53, Rea56, LLM59]. Structures [Sno57]. Studies [Bec51, CFJ56, CHK59, De 52, FKR53, Kap55, Sob56]. Study [Clo53, Kay53]. Subject [GM55b]. Submarine [EH57, Gil51b, LFB57, Zaj57]. Submerged [Gil51b]. Substituted [SD59]. Successive [Des57]. Sufficient [Ben57b]. Suitable [Bob57]. Super [Kay53]. Super-High [Kay53]. Supply [Blo53]. Suppression [Rut58]. Surface [Bac58, BS56b, BB53, BG56, FM56]. Surfaces [ATS59, GB56, Hop52]. Sweeping [Hin50]. Swept [Pri53, Rea56]. Switch [Des57, Ell55, Hay52]. Switches [HP55, ME55a]. Switching [Ada56, And52, Ata53b, Ata53a, Ata54, Ata55b, Ata55a, BH55, BHM56, CO52, Clo53, EM55, FG58, FR58, Fre58, GGI58, Han56, Joe56a, Joe58, Kel52, Lee55, MV52, MW52, Mye52, Raf54, Sch51, SY58, Sem58, Shi52, Tow57, WEP55, YCU58, Gil51a, HS55, LLM59, Lee59, Sta52]. Symbols [Ano54q]. symmetry [McC56a]. Synthesis [Dar51a, Kin59a, NM59, Sem58, Dar52]. Synthesizing [Mea55]. Synthetic [Kin59b]. System [Ada56, Ano50f, Ano50g, Ano50h, Ano51k, Ano53m, BHW59, BG57, BHP57, Car51, DK55, DKK53, Ear55, EM52, EEKG53, FG58, FW51, Gal59, GTW53, HB57, HE59, HHH59, HRR52, JLL57, Joe58, KP51, KF53, Ket59, Kin54, KM56, LFB57, Lev58, LTLF57, LMM55, MV52, MMVH57, MLD53, MHEG57, RG53, RSF51, SY58, TK57, WEP55, YCU58, LLM59, Ano50a, Ano50f, Ano50g, Ano50h, Ano51a, Ano51b, Ano51c, Ano51k, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano52f, Ano53a, Ano53b, Ano53c, Ano53d, Ano53e, Ano53f, Ano54a, Ano54b, Ano54c, Ano54d, Ano54e, Ano54f, Ano54n, Ano54o, Ano54p, Ano55a, Ano55b, Ano55c, Ano55d, Ano55e, Ano55f, Ano55p, Ano55q, Ano55r, Ano55s, Ano55t, Ano55u, Ano56a, Ano56b, Ano56c, Ano56d]. System [Ano56e, Ano56f, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano57f, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano57s, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano59m, Sha51b, Sha51e, Sha51c, Sha51d]. Systems [Ata55b, Bab53, BCR55, Ble56, BD53, EH57, Fre58, Har59, Joe56a, Kel52, MA54, Mye52, Raf54, Rut58, Sch51, Shi52, Tow57, Uhl55].

T [Ano57t]. Table [Bea58]. Tables [SH58, Tho56]. Tamping [WM59]. Tandem [Ada56]. Taper [Ung58a]. Tapered [Coo55, Lou55]. TASI [BF59]. Tchebycheff [Dar52]. TD [BG57, RSF51, WEP55]. TD-2 [BG57, RSF51, WEP55]. TE01 [Mar58]. Technical [Ano50a, Ano50f, Ano50g, Ano50h, Ano51a, Ano51b, Ano51c, Ano51k, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano52f, Ano53a, Ano53b, Ano53c, Ano53d, Ano53e, Ano53f, Ano54a, Ano54b, Ano54c, Ano54d, Ano54e, Ano54f, Ano54n, Ano54o, Ano54p, Ano55a, Ano55b, Ano55c, Ano55d, Ano55e, Ano55f, Ano55p, Ano55q, Ano55r, Ano55s, Ano55t, Ano55u, Ano56a, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano57f, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano57s, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano59m]. Techniques [Ear55, Feh57, Joe58, MW52, Tru54, Sob56]. Technology [Mar51]. Telegraph [HRR52]. Telegraphist [Sch52, Sch55]. Telephone [Ano51j, Ata53b, Ata53a, Ata54, Ata55b, Ata55a, Ben57a, Ben57b, Ben53a, BCH55, BHP55, BH55, Car51, CHS54, CO52, CW52, CR59, FKR53, Gil51b, GM53, Gre58, Hay52, HV55, Hun53, IT51, Joe56a, KP51, Kel52, Mar51, MW52, MRS54, MM57, MR50, MM51, MHEG57, Mye52, Pfe54, Pil52, Raf54, Rio51, Rob56, Sch51, Sha50, Sha51b, Sha51e, Sha51c, Sha51d, Tow57, Wag54a, Win53]. Television [Goo51, Har52, Jen51, Jul59, Kre52, RG53, RL55]. Temperature [Mas51]. Temperatures [Kin59b, Mas51, Pee51]. Ten [Bak58]. Tensor [Mas50]. term [Ben53b]. terminal [Gil51a]. Terminals [RG53]. Termination [Jac58]. Test [BD53, Col53a]. Testing [CS56, FPS56, Han56, OW59, Rob56, SG59]. Tests [Col53a, Rob58, Van53]. Tetrode [LS56]. Their [Gra53, MA54]. Theoretical [Sun54a, Sun54b]. Theories [Wil56]. Theory [Ata53b, AC54, Bar50, Ear53, Fuc59, McM52a, McM52b, Mer51, Mil54a, Mor52b, Mor52a, Mor57, Pee54b, Rea56, Ric50b, Suh53, SW54c, Tie56, Ung58b, Van50, FCH57]. Thermally [ATS59]. Thermomagnetic [HGK59]. Thickness [Epp54, Woj54]. Thicknesses [Fuc59]. Thin [BS56b, Hei51]. Third [Pie50d]. Thirtieth [Ano52m]. Three [Daw57, DSDS59, RM58, SDSD59]. Three-Conductor [RM58]. Three-Level [DSDS59, SDSD59]. Throwdown [FKR53]. Time [Log54, Pee54b, Rio51, Sho54, Vau59, Bur59, Sob56]. Time-Separation [Vau59]. times [Ben59]. Timing [DP58, De 58, Row58, Sun57]. Titanate [MS58]. Tolerance [Eva58]. Toll [BHM56, Pil52, Shi52, Wil56]. Tone [MPW58]. Tools [Mal53]. Topics [SW54a, SW54b, SW54c]. total [McC56a]. Traffic [Ben57a, FKR53, Hay52, Rio51, Tru54, Wil56]. Transatlantic [JLL57, KR57, LFB57, MMVH57, MHEG57]. transcendental [Zin52]. Transcribed [Men52]. Transfer [DB54]. Transformation [Hel59]. Transformer [Lev58]. Transformers [Ear55]. Transient [Ell55]. Transistor [Ano56m, Bak58, Ban54, Ble56, Blo53, DR55, Ear54, FM56, Lee56, LS56, Mor52c, NA59, Ten56, Tho59, WR51]. Transistorized [BHD59, Wra56]. Transistors [And52, CHS54, Ear53, EM55, ME55b, TT56, WP51]. Transit [Sho54]. Transition [Mor58]. Translator [BT54, BHM56, HN53]. Translators [Sch51]. Transmission [Ben58, BCH55, Car51, DK55, FPS56, FW51, GM55a, Hei51, HV55, Hun53, Jac58, KM56, Kin59c, MM57, Mor52b, Mor52a, RM58, Rai59, Ric59, RL55, Sou50, Sun54a, Sun54b, Sun59, TD53, Vaa53, War58, You52, Ric58, Ung57a, Ung57b]. Transport [Her55]. Transpositions [RM58]. Transversal [BG57]. Transverse [PW56, SW54b]. Traveling [Cut56, DSDS59, LMM56, Pie50a, Pie50c, Pie50d, Pie50b, Tie56]. Traveling-Wave [DSDS59, LMM56, Pie50a, Pie50c, Pie50d, Pie50b]. Treatments [FM56]. Triode [BM50, MR50]. Triodes [Ear54]. Tropospheric [CHK59]. Trunk [FL56, Joh54, Tru54]. Trunks [FPS56, Hun53, Ben59]. Tube [Aul59, BDS58, Jen51, LMM56]. Tubes [Cut56, FW51, HCM55, MMVH57, Pie50a, Pie50c, Pie50d, Pie50b, Pie54, Stu53, TD53, Tow57]. Tunable [Ree55]. Twin [AC54]. Twistor [Bob57]. Two [Ric50a]. Type [Car51, Col53a, DRS56, EM52, HGK59, Sha56, Vaa53]. Type-O [EM52].

U.S.A. [Wil56]. UHF [Rig57a]. Ultrasonic [MM52, Mas55]. Uniformity [BS56a]. Unit [Tho56]. Up-Converter [Lee58]. Use [Bec51, Blo53, BS56b, Clo51, Ell59, FG58, SDSD59]. Used [Sch51]. Using [Jen51, Joe58, Kis58, MM57, Dar52].

V [Ata55a]. Valley [Her55]. Valve [Rau59, Sch57]. Valve-Controlled [Rau59]. Vapors [GS57]. Variable [GM59, Lee58]. Variable-Capacitance [Lee58]. Variable-Length [GM59]. Variations [GM55b]. Velocities [PW56]. Velocity [Bur58, Coo55]. Verification [LLM59]. Vernier [Kro57]. Very [De 52, Mas51]. Video [DK55]. View [BT58a, BT58b]. Visual [BHW59]. Voice [MM57]. Voltage [BG56, ME55a, Ree55, VP57]. Volume [CR59].

W [Bec51]. W. [Ano57t]. Wafer [Sha56]. Wafer-Type [Sha56]. Wall [Fuc59]. Walls [Gal54, Gal55]. Warped [Fox55]. Wave [Cut56, DSDS59, Fox55, Kin56, KH51, LMM56, Mar58, Mas55, Mil54a, Pie50a, Pie50c, Pie50d, Pie50b, Pie54, RL54, RP59, Rig59, Sha56, SW54a, SW54b, SW54c, Tie56, Ung57a, Ung57b]. Waveguide [Bec56, Fuc59, Lec59, Mar58, Mil54a, Mil54b, MY56, Sei57, SF59, Sou50, Ung58a, Ung58b, War58, Mor57, Ung57a]. Waveguides [Ros58, Sch52]. Wavelengths [CH56]. Waves [BH52, Pie51b, PW56, Ric54, Ung57c, War58, Har50b, Hol51]. Wear [MW52]. Welding [Qui54, SB59, Sum54]. Wide [Hin50]. Windings [Log54, Pee51]. Wire [BRT56, BCS54, GM53, Qui54]. Within [Mar58]. Wood [Col53a, Col53b, Egg50]. Work [Dar51b]. Worked [Hei51]. Working [Wil53]. Wrapped [Ell59, Mal53, MO53, MA54, McR53, Van53]. Writings [Ano51i].

Years [Bak58, Mar51]. Yield [Eva58].

Zero [Mas51]. Zeros [Bea58]. Zone [BS56a, Rob58].

References

Atalla:1954:TOC

[AC54] M. M. Atalla and Miss R. E. Cox. Theory of open-contact performance of twin contacts. The Bell System Technical Journal, 33(6):1373–1386, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                                 tronic). URL http://bstj.bell-
                                                 labs.com/BSTJ/images/Vol33/
                                                 bstj33-6-1373.pdf;   http://
                                                 www.alcatel-lucent.com/bstj/
                                                 vol33-1954/articles/bstj33-6-
                                                 1373.pdf.

Adam:1956:CTL

[Ada56] A. O. Adam. Crossbar tandem as a long distance switching system. The Bell System Technical Journal, 35(1):91–108, January 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol35/bstj35-
1-91.pdf;   http://www.alcatel-
lucent.com/bstj/vol35-1956/
articles/bstj35-1-91.pdf.

Alley:1953:RNM

[All53] R. E. Alley, Jr. A review of new magnetic phenomena. The Bell System Technical Journal, 32 (5):1155–1172, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-5-1155.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-5-
1155.pdf.

Anderson:1952:TSC

[And52] A. Eugene Anderson. Transistors in switching circuits. The Bell System Technical Journal, 31(6):1207–1249, November 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-6-1207.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-6-
1207.pdf.

Anonymous:1950:ATA

[Ano50a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 29(1):142–145, January 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-1-142.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-1-
                                        142.pdf.

Anonymous:1950:CIa

[Ano50b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 29(1):146, January 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-1-146.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-1-
                                        146.pdf.

Anonymous:1950:CIb

[Ano50c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 29(2):294, April 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-2-294.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-2-
                                        294.pdf.

Anonymous:1950:CIc

[Ano50d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 29(3):468, July 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-3-468.pdf;   http://

www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-3-
468.pdf.

Anonymous:1950:CId

[Ano50e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 29(4):674–675, October 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-4-674.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-4-
674.pdf.

Anonymous:1950:TAB

[Ano50f] Anonymous. Technical articles by Bell System authors not appearing in the Bell System Technical Journal. The Bell System Technical Journal, 29(2):287–293, April 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-2-287.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-2-
287.pdf.

Anonymous:1950:TPBa

[Ano50g] Anonymous. Technical publications by Bell System authors other than in the Bell System Technical Journal. The Bell System Technical Journal, 29(3):461–467, July 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-3-461.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-3-
                                        461.pdf.

Anonymous:1950:TPBb

[Ano50h] Anonymous. Technical publications by Bell System authors other than in the Bell System Technical Journal. The Bell System Technical Journal, 29(4):672–673, October 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-4-672.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-4-
                                        672.pdf.

Anonymous:1951:ABSa

[Ano51a] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 30 (2):473–487, April 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-2-473.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-2-
                                        473.pdf.

Anonymous:1951:ABSb

[Ano51b] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 30(3):765–776, July 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol30/
bstj30-3-765.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-3-
765.pdf.

Anonymous:1951:ABSc

[Ano51c] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 30(4): 1244–1253, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1244.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1244.pdf.

Anonymous:1951:CIa

[Ano51d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 30(1):211–212, January 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-1-211.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-1-
211.pdf.

Anonymous:1951:CIb

[Ano51e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 30(2):488–490, April 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-2-488.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-2-
                                        488.pdf.

Anonymous:1951:CIc

[Ano51f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 30(3):777–778, July 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-3-777.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-3-
                                        777.pdf.

Anonymous:1951:CId

[Ano51g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 30(4):1038–1040, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-4-1038.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-4-
                                        1038.pdf.

Anonymous:1951:CIe

[Ano51h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 30(4):1254–1255, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-4-1254.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-4-
                                        1254.pdf.

Anonymous:1951:PWC

[Ano51i] Anonymous. Published writings of C. J. Davisson. The Bell System Technical Journal, 30(4): 1035–1037, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1035.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1035.pdf.

Anonymous:1951:SFA

[Ano51j] Anonymous. The seventy-fifth anniversary of the telephone. The Bell System Technical Journal, 30 (2):213–214, April 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-2-213.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-2-
213.pdf.

Anonymous:1951:TPB

[Ano51k] Anonymous. Technical publications by Bell System authors other than in the Bell System Technical Journal. The Bell System Technical Journal, 30(1):205–210, January 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-1-205.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-1-
205.pdf.

Anonymous:1952:ABSa

[Ano52a] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 31(1): 201–212, January 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-1-201.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-1-
                                        201.pdf.

Anonymous:1952:ABSb

[Ano52b] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 31 (2):405–408, March 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-2-405.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-2-
                                        405.pdf.

Anonymous:1952:ABSc

[Ano52c] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 31(3):601–607, May 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-3-601.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol31-1952/articles/bstj31-3-
601.pdf.

Anonymous:1952:ABSd

[Ano52d] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 31(4):816–819, July 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-4-816.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
816.pdf.

Anonymous:1952:ABSe

[Ano52e] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 31(5): 1013–1019, September 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-1013.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
1013.pdf.

Anonymous:1952:ABSf

[Ano52f] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 31(6): 1250–1255, November 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/

                                        bstj31-6-1250.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-6-
                                        1250.pdf.

Anonymous:1952:CIa

[Ano52g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 31(1):213–215, January 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-1-213.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-1-
                                        213.pdf.

Anonymous:1952:CIb

[Ano52h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 31(2):409–410, March 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-2-409.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-2-
                                        409.pdf.

Anonymous:1952:CIc

[Ano52i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 31(3):608–609, May 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-3-608.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-3-
                                        608.pdf.

Anonymous:1952:CId

[Ano52j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 31(4):820–822, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-4-820.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
820.pdf.

Anonymous:1952:CIe

[Ano52k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 31(5):1020–1022, September 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-1020.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
1020.pdf.

Anonymous:1952:CIf

[Ano52l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 31(6):1256–1257, November 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-6-1256.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-6-
1256.pdf.

Anonymous:1952:TA

[Ano52m] Anonymous. Thirtieth anniversary. The Bell System Technical Journal, 31(4):611–612, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-4-611.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
611.pdf.

Anonymous:1953:ABSa

[Ano53a] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 32(1): 255–260, January 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-1-255.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-1-
255.pdf.

Anonymous:1953:ABSb

[Ano53b] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 32 (2):506–518, March 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-2-506.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-2-
506.pdf.

Anonymous:1953:ABSc

[Ano53c] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 32(3):767–774, May 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-3-767.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-3-
767.pdf.

Anonymous:1953:ABSd

[Ano53d] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 32(4): 1007–1014, July 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-4-1007.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-4-
1007.pdf.

Anonymous:1953:ABSe

[Ano53e] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 32(5): 1257–1265, September 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-5-1257.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-5-
1257.pdf.

Anonymous:1953:ABSf

[Ano53f] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 32(6): 1507–1513, November 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-6-1507.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-6-
                                        1507.pdf.

Anonymous:1953:CIa

[Ano53g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 32(1):261–263, January 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-1-261.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-1-
                                        261.pdf.

Anonymous:1953:CIb

[Ano53h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 32(2):519–521, March 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-2-519.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-2-
                                        519.pdf.

Anonymous:1953:CIc

[Ano53i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 32(3):775–777, May 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-3-775.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-3-
775.pdf.

Anonymous:1953:CId

[Ano53j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 32(4):1015–1018, July 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-4-1015.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-4-
1015.pdf.

Anonymous:1953:CIe

[Ano53k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 32(5):1267–1270, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-5-1267.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-5-
1267.pdf.

Anonymous:1953:CIf

[Ano53l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 32(6):1515–1517, November 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/

                                        bstj32-6-1515.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-6-
                                        1515.pdf.

Anonymous:1953:LCS

[Ano53m] Anonymous. The L3 coaxial system — foreword. The Bell System Technical Journal, 32(4):779–780, July 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-4-779.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-4-
                                        779.pdf.

Anonymous:1954:ABSa

[Ano54a] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 33(1): 260–273, January 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-1-260.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-1-
                                        260.pdf.

Anonymous:1954:ABSb

[Ano54b] Anonymous. Abstracts of Bell System technical papers not published in this journal. The Bell System Technical Journal, 33 (2):505–513, March 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/

bstj33-2-505.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-2-
505.pdf.

Anonymous:1954:BSTa

[Ano54c] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 33(3):789–796, May 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-3-789.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-3-
789.pdf.

Anonymous:1954:BSTb

[Ano54d] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 33(4):1011–1016, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-1011.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
1011.pdf.

Anonymous:1954:BSTc

[Ano54e] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 33(5):1195–1200, September 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-5-1195.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-5-
                                        1195.pdf.

Anonymous:1954:BSTd

[Ano54f] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 33(6):1387–1391, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-6-1387.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-6-
                                        1387.pdf.

Anonymous:1954:CIa

[Ano54g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 33(1):274–275, January 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-1-274.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-1-
                                        274.pdf.

Anonymous:1954:CIb

[Ano54h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 33(2):514–516, March 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-2-514.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-2-
                                        514.pdf.

Anonymous:1954:CIc

[Ano54i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 33(3):797–798, May 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-3-797.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-3-
797.pdf.

Anonymous:1954:CId

[Ano54j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 33(4):1019–1022, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-1019.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
1019.pdf.

Anonymous:1954:CIe

[Ano54k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 33(5):1206–1208, September 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-5-1206.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-5-
1206.pdf.

Anonymous:1954:CIf

[Ano54l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 33(6):1398–1400, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-6-1398.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-6-
1398.pdf.

Anonymous:1954:DRI

[Ano54m] Anonymous. Design of relays — introduction. The Bell System Technical Journal, 33(1):1–2, January 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol33-
1954/articles/bstj33-1-1.pdf.

Anonymous:1954:RMBa

[Ano54n] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 33(4):1017–1018, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-1017.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
1017.pdf.

Anonymous:1954:RMBb

[Ano54o] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 33(5):1201–1205, September 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-5-1201.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-5-
1201.pdf.

Anonymous:1954:RMBc

[Ano54p] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 33(6):1392–1397, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-6-1392.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-6-
1392.pdf.

Anonymous:1954:S

[Ano54q] Anonymous. Symbols. The Bell System Technical Journal, 33(1): 257–259, January 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-1-257.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-1-
257.pdf.

Anonymous:1955:BSTa

[Ano55a] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(1):221–229, January 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-1-221.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-1-
                                        221.pdf.

Anonymous:1955:BSTb

[Ano55b] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(2):443–447, March 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-2-443.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-2-
                                        443.pdf.

Anonymous:1955:BSTc

[Ano55c] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(3):665–670, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-3-665.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-3-
                                        665.pdf.

Anonymous:1955:BSTd

[Ano55d] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(4):871–875, July 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-4-871.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-4-
871.pdf.

Anonymous:1955:BSTe

[Ano55e] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(5):1103–1108, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-5-1103.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-5-
1103.pdf.

Anonymous:1955:BSTf

[Ano55f] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(6):1317–1322, November 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-6-1317.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-6-
1317.pdf.

Anonymous:1955:CIa

[Ano55g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 34(1):234–236, January 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/

                                        bstj34-1-234.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-1-
                                        234.pdf.

Anonymous:1955:CIb

[Ano55h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 34(2):450–452, March 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-2-450.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-2-
                                        450.pdf.

Anonymous:1955:CIc

[Ano55i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 34(3):672–676, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-3-672.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-3-
                                        672.pdf.

Anonymous:1955:CId

[Ano55j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 34(4):879–882, July 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-4-879.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-4-
                                        879.pdf.

Anonymous:1955:CIe

[Ano55k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 34(5):1112–1114, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-5-1112.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-5-
1112.pdf.

Anonymous:1955:CIf

[Ano55l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 34(6):1327–1330, November 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-6-1327.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-6-
1327.pdf.

Anonymous:1955:Ca

[Ano55m] Anonymous. Correction. The Bell System Technical Journal, 34(1):104, January 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-1-104.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-1-
104.pdf.

Anonymous:1955:Cb

[Ano55n] Anonymous. Correction. The Bell System Technical Journal, 34 (5):994, September 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-5-994.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-5-
994.pdf.

Anonymous:1955:GC

[Ano55o] Anonymous. George A. Campbell, 1870 — 1954. The Bell System Technical Journal, 34 (1):1–4, January 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol34-
1955/articles/bstj34-1-1.pdf.

Anonymous:1955:RMBa

[Ano55p] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(1):230–233, January 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-1-230.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-1-
230.pdf.

Anonymous:1955:RMBb

[Ano55q] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(2):448–449, March 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-2-448.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-2-
448.pdf.

Anonymous:1955:RMBc

[Ano55r] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(3):670–671, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-3-670.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-3-
670.pdf.

Anonymous:1955:RMBd

[Ano55s] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(4):876–878, July 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-4-876.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-4-
876.pdf.

Anonymous:1955:RMBe

[Ano55t] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(5):1109–1111, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-5-1109.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-5-
                                        1109.pdf.

Anonymous:1955:RMBf

[Ano55u] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 34(6):1323–1326, November 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-6-1323.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-6-
                                        1323.pdf.

Anonymous:1956:BSTa

[Ano56a] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(1):235–241, January 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-1-235.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-1-
                                        235.pdf.

Anonymous:1956:BSTb

[Ano56b] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(2):519–526, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-2-519.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-2-
519.pdf.

Anonymous:1956:BSTc

[Ano56c] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(3):751–758, May 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-3-751.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-3-
751.pdf.

Anonymous:1956:BSTd

[Ano56d] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(4):973–978, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-4-973.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-4-
973.pdf.

Anonymous:1956:BSTe

[Ano56e] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(5):1223–1229, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-1223.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1223.pdf.

Anonymous:1956:BSTf

[Ano56f] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(6):1454–1460, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-6-1454.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-6-
                                        1454.pdf.

Anonymous:1956:CIa

[Ano56g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 35(1):244–247, January 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-1-244.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-1-
                                        244.pdf.

Anonymous:1956:CIb

[Ano56h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 35(2):531–534, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-2-531.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-2-
531.pdf.

Anonymous:1956:CIc

[Ano56i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 35(3):762–765, May 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-3-762.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-3-
762.pdf.

Anonymous:1956:CId

[Ano56j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 35(4):985–989, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-4-985.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-4-
985.pdf.

Anonymous:1956:CIe

[Ano56k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 35(5):1233–1237, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-5-1233.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1233.pdf.

Anonymous:1956:CIf

[Ano56l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 35(6):1465–1466, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-6-1465.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-6-
                                        1465.pdf.

Anonymous:1956:NPP

[Ano56m] Anonymous. Nobel Prize in Physics awarded to transistor inventors. The Bell System Technical Journal, 35(6):i–iv, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-6-i.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol35-
                                        1956/articles/bstj35-6-i.pdf.

Anonymous:1956:RMBa

[Ano56n] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(1):242–243, January 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-1-242.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol35-1956/articles/bstj35-1-
242.pdf.

Anonymous:1956:RMBb

[Ano56o] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(2):527–530, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-2-527.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-2-
527.pdf.

Anonymous:1956:RMBc

[Ano56p] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(3):759–761, May 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-3-759.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-3-
759.pdf.

Anonymous:1956:RMBd

[Ano56q] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(4):979–984, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/

                                        bstj35-4-979.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-4-
                                        979.pdf.

Anonymous:1956:RMBe

[Ano56r] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(5):1230–1232, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-1230.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1230.pdf.

Anonymous:1956:RMBf

[Ano56s] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 35(6):1461–1464, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-6-1461.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-6-
                                        1461.pdf.

Anonymous:1957:BSTa

[Ano57a] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(1):327–334, January 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol36/
bstj36-1-327.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-1-
327.pdf.

Anonymous:1957:BSTb

[Ano57b] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(2):577–582, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-2-577.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-2-
577.pdf.

Anonymous:1957:BSTc

[Ano57c] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(3):813–822, May 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-3-813.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-3-
813.pdf.

Anonymous:1957:BSTd

[Ano57d] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(4):1035–1042, July 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/

                                        bstj36-4-1035.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-4-
                                        1035.pdf.

Anonymous:1957:BSTe

[Ano57e] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(5):1308–1313, September 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-5-1308.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-5-
                                        1308.pdf.

Anonymous:1957:BSTf

[Ano57f] Anonymous. Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(6):1501–1507, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-6-1501.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-6-
                                        1501.pdf.

Anonymous:1957:CIa

[Ano57g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 36(1):338–348, January 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-1-338.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol36-1957/articles/bstj36-1-
338.pdf.

Anonymous:1957:CIb

[Ano57h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 36(2):588–591, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-2-588.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-2-
588.pdf.

Anonymous:1957:CIc

[Ano57i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 36(3):828–830, May 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-3-828.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-3-
828.pdf.

Anonymous:1957:CId

[Ano57j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 36(4):1045–1046, July 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-4-1045.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-4-
1045.pdf.

Anonymous:1957:CIe

[Ano57k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 36(5):1317–1318, September 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-5-1317.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-5-
                                        1317.pdf.

Anonymous:1957:CIf

[Ano57l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 36(6):1511–1513, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-6-1511.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-6-
                                        1511.pdf.

Anonymous:1957:E

[Ano57m] Anonymous.  Errata.  The

Bell System Technical Journal, 36(2):496, March 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-2-496.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-2-

496.pdf. See [FM56] and [LFB57].

Anonymous:1957:RMBa

[Ano57n] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(1):335–337, January 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-1-335.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-1-
335.pdf.

Anonymous:1957:RMBb

[Ano57o] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(2):583–587, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-2-583.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-2-
583.pdf.

Anonymous:1957:RMBc

[Ano57p] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(3):823–827, May 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-3-823.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-3-
823.pdf.

Anonymous:1957:RMBd

[Ano57q] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(4):1043–1044, July 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-4-1043.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-4-
                                        1043.pdf.

Anonymous:1957:RMBe

[Ano57r] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(5):1313–1317, September 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-5-1313.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-5-
                                        1313.pdf.

Anonymous:1957:RMBf

[Ano57s] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 36(6):1508–1510, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-6-1508.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol36-1957/articles/bstj36-6-
1508.pdf.

Anonymous:1957:WDB

[Ano57t] Anonymous. W. D. Bulloch appointed editor of B. S. T. J. The Bell System Technical Journal, 36(3):710, May 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-3-710.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-3-
710.pdf.

Anonymous:1958:CIa

[Ano58a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 37(1):285–288, January 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-1-285.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-1-
285.pdf.

Anonymous:1958:CIb

[Ano58b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 37(2):576–579, March 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-2-576.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-2-
576.pdf.

Anonymous:1958:CIc

[Ano58c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 37(3):831–834, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-3-831.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-3-
                                        831.pdf.

Anonymous:1958:CId

[Ano58d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 37(4):1088–1090, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-4-1088.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-4-
                                        1088.pdf.

Anonymous:1958:CIe

[Ano58e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 37(5):1323–1326, September 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-5-1323.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-5-
                                        1323.pdf.

Anonymous:1958:CIf Journal, 37(6):1670–1672, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-6-1670.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-6-
1670.pdf.

Anonymous:1958:RMBa

[Ano58g] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 37(1):283–285, January 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-1-283.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-1-
283.pdf.

Anonymous:1958:RMBb

[Ano58h] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 37(2):571–575, March 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-2-571.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-2-
571.pdf.

Anonymous:1958:RMBc

[Ano58i] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 37(3):827–830, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-827.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-3-
827.pdf.

Anonymous:1958:RMBd

[Ano58j] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 37(4):1085–1087, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-4-1085.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-4-
1085.pdf.

Anonymous:1958:RMBe

[Ano58k] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 37(5):1319–1322, September 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-5-1319.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-5-
1319.pdf.

Anonymous:1958:RMBf

[Ano58l] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 37(6):1663–1669, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-6-1663.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-6-
1663.pdf.

Anonymous:1959:CIa

[Ano59a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 38(1):300–304, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-1-300.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-1-
300.pdf.

Anonymous:1959:CIb

[Ano59b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 38(2):606–610, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-606.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
606.pdf.

Anonymous:1959:CIc

[Ano59c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 38(3):905–907, May 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-3-905.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-3-
                                        905.pdf.

Anonymous:1959:CId

[Ano59d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 38(4):1065–1066, July 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-4-1065.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-4-
                                        1065.pdf.

Anonymous:1959:CIe

[Ano59e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 38(5):1335–1336, September 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-5-1335.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-5-
                                        1335.pdf.

Anonymous:1959:CIf Journal, 38(6):1601–1603, November 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1601.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1601.pdf.

Anonymous:1959:HRN

[Ano59g] Anonymous. H. S. Renne is new editor of B.S.T.J. The Bell System Technical Journal, 38(1):18, January 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol38/bstj38-
1-18.pdf;   http://www.alcatel-
lucent.com/bstj/vol38-1959/
articles/bstj38-1-18.pdf.

Anonymous:1959:RMBa

[Ano59h] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 38(1):297–299, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-1-297.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-1-
297.pdf.

Anonymous:1959:RMBb

[Ano59i] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 38(2):603–605, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-2-603.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-2-
                                        603.pdf.

Anonymous:1959:RMBc

[Ano59j] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 38(3):899–904, May 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-3-899.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-3-
                                        899.pdf.

Anonymous:1959:RMBd

[Ano59k] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 38(4):1061–1064, July 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-4-1061.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-4-
                                        1061.pdf.

Anonymous:1959:RMBe

[Ano59l] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 38(5):1331–1334, September 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-5-1331.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-5-
1331.pdf.

Anonymous:1959:RMBf

[Ano59m] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 38(6):1595–1600, November 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1595.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1595.pdf.

Atalla:1953:AECb

[Ata53a] M. M. Atalla. Arcing of electrical contacts in telephone switching circuits — Part II — characteristics of the short arc. The Bell System Technical Journal, 32(6):1493–1506, November 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-6-1493.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-6-
1493.pdf.

Atalla:1953:AECa

[Ata53b] M. M. Atalla. Arcing of electrical contacts in telephone switching circuits: Part I — theory of the initiation of the short arc. The Bell System Technical Journal, 32(5):1231–1244, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-5-1231.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-5-
                                        1231.pdf.

Atalla:1954:AEC

[Ata54] M. M. Atalla. Arcing of electrical contacts in telephone switching circuits: Part III — discharge phenomena on break of inductive circuits. The Bell System Technical Journal, 33(3):535–558, May 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-3-535.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-3-
                                        535.pdf.

Atalla:1955:AECb

[Ata55a] M. M. Atalla. Arcing of electrical contacts in telephone switching circuits: Part V — mechanisms of the short arc and erosion of contacts. The Bell System Technical Journal, 34(5):1081–1102, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/

bstj34-5-1081.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-5-
1081.pdf.

Atalla:1955:AECa

[Ata55b] M. M. Atalla. Arcing of electrical contacts in telephone switching systems: Part IV — mechanism of the initiation of the short arc. The Bell System Technical Journal, 34(1):203–220, January 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-1-203.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-1-
203.pdf.

Atalla:1959:SSS

[ATS59] M. M. Atalla, E. Tannenbaum, and E. J. Scheibner. Stabilization of silicon surfaces by thermally grown oxides. The Bell System Technical Journal, 38(3):749–783, May 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-3-749.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-3-
749.pdf.

Ault:1959:SHS

[Aul59] C. F. Ault. Stable high-speed digital-to-analog conversion for storage tube deflection. The Bell System Technical Journal, 38 (2):445–465, March 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-2-445.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-2-
                                        445.pdf.

Babcock:1953:IIR

[Bab53] Wallace C. Babcock. Intermodulation interference in radio systems: Frequency of occurrence and control by channel selection. The Bell System Technical Journal, 32 (1):63–73, January 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol32/bstj32-
                                        1-63.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol32-1953/
                                        articles/bstj32-1-63.pdf.

Backenstoss:1958:ESC

[Bac58] Gerhard Backenstoss. Evaluation of the surface concentration of diffused layers in silicon. The Bell System Technical Journal, 37(3):699–710, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-3-699.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-3-
                                        699.pdf.

Baker:1958:FTY

[Bak58] W. O. Baker. The first ten years of the transistor: A guest editorial. The Bell System Technical Journal, 37(5):i–vi, September 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-5-i.pdf;   http://www.
alcatel-lucent.com/bstj/vol37-
1958/articles/bstj37-5-i.pdf.

Bangert:1954:TNE

[Ban54] J. T. Bangert. The transistor as a network element. The Bell System Technical Journal, 33 (2):329–352, March 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-2-329.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-2-
329.pdf.

Bardeen:1950:TRB

[Bar50] J. Bardeen. Theory of relation between hole concentration and characteristics of germanium point contacts. The Bell System Technical Journal, 29(4):469–495, October 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-4-469.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-4-
469.pdf.

Brattain:1953:SPG

[BB53] Walter H. Brattain and John Bardeen. Surface properties of germanium. The Bell System Technical Journal, 32(1):1–41, January 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol32-
                                        1953/articles/bstj32-1-1.pdf.

Blye:1955:RTT

[BCH55] P. W. Blye, O. H. Coolidge, and H. R. Huntley. A revised telephone transmission rating plan. The Bell System Technical Journal, 34(3):453–472, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-3-453.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-3-
                                        453.pdf.

Bennett:1955:IIF

[BCR55] W. R. Bennett, H. E. Curtis, and S. O. Rice. Interchannel interference in FM and PM systems under noise loading conditions. The Bell System Technical Journal, 34(3):601–636, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-3-601.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-3-
                                        601.pdf.

Brunner:1954:WSM

[BCS54] A. J. Brunner, H. E. Cosson, and R. W. Strickland. Wire straightening and molding for wire spring relays. The Bell System Technical Journal, 33(4):859–884, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-859.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
859.pdf.

Burns:1953:ALI

[BD53] R. W. Burns and J. W. Dehn.

Automatic line insulation test equipment for local crossbar systems. The Bell System Technical Journal, 32(3):627–646, May 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-3-627.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-3-
627.pdf.

Brangaccio:1958:BOT

[BDS58] D. J. Brangaccio, A. F. Dietrich, and J. W. Sullivan. Broadband oscilloscope tube. The Bell System Technical Journal, 37 (2):447–460, March 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-2-447.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-2-
447.pdf.

Beattie:1958:TFZ

[Bea58] Curtis L. Beattie. Table of first 700 zeros of Bessel functions — J_{l}(x) and J′_{l}(x). The Bell System Technical Journal, 37(3):689–697, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-3-689.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-3-
                                        689.pdf.

Becker:1951:UFE

[Bec51] J. A. Becker. The use of the field emission electron microscope in absorption studies of W on W and Ba on W. The Bell System Technical Journal, 30(4):907–932, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-4-907.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-4-
                                        907.pdf.

Beck:1956:WIM

[Bec56] A. C. Beck. Waveguide investigations with millimicrosecond pulses. The Bell System Technical Journal, 35(1):35–65, January 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol35/bstj35-
                                        1-35.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol35-1956/
                                        articles/bstj35-1-35.pdf.

Bennett:1953:ICT

[Ben53a] A. F. Bennett. An improved circuit for the telephone set. The Bell System Technical Journal, 32(3):611–626, May 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-3-611.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-3-
611.pdf.

Bennett:1953:CMC

[Ben53b] W. R. Bennett. The correlatograph. A machine for continuous display of short term correlation. The Bell System Technical Journal, 32(5):1173–1185, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-5-1173.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-5-
1173.pdf.

Benes:1957:FTT

[Ben57a] V. E. Beneš. Fluctuation of telephone traffic. The Bell System Technical Journal, 36 (4):965–973, July 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-4-965.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-4-
965.pdf.

Benes:1957:SSS

[Ben57b] V. E. Beneš. A sufficient set of statistics for a simple telephone exchange model. The Bell System Technical Journal, 36(4):939–964, July 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-4-939.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-4-
                                        939.pdf.

Bennett:1958:SRD

[Ben58] W. R. Bennett. Statistics of regenerative digital transmission. The Bell System Technical Journal, 37(6):1501–1542, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-6-1501.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-6-
                                        1501.pdf.

Benes:1959:TNE

[Ben59] Václav E. Beneš. On trunks with negative exponential holding times serving a renewal process. The Bell System Technical Journal, 38(1):211–258, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-1-211.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-1-
                                        211.pdf.

Bullington:1959:EAT

[BF59] K. Bullington and J. M. Fraser. Engineering aspects of TASI. The Bell System Technical Journal, 38 (2):353–364, March 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-353.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
353.pdf.

Brattain:1955:EIB

[BG55] W. H. Brattain and C. G. B. Garrett. Experiments on the interface between germanium and an electrolyte. The Bell System Technical Journal, 34(1):129–176, January 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-1-129.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-1-
129.pdf.

Brattain:1956:CMF

[BG56] W. H. Brattain and C. G. B.

Garrett. Combined measurements of field effect, surface photovoltage and photoconductivity. The Bell System Technical Journal, 35(5):1019–1040, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-5-1019.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-5-
1019.pdf.

Bellows:1957:ETE

[BG57] B. C. Bellows and R. S. Graham. Experimental transversal equalizer for TD-2 radio relay system. The Bell System Technical Journal, 36(6):1429–1450, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-6-1429.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-6-
                                        1429.pdf.

Baker:1952:IPM

[BH52] W. O. Baker and J. H. Heiss. Interaction of polymers and mechanical waves. The Bell System Technical Journal, 31(2):306–356, March 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-2-306.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-2-
                                        306.pdf.

Brewer:1955:TSN

[BH55] S. T. Brewer and G. Hecht. A telephone switching network and its electronic controls. The Bell System Technical Journal, 34 (2):361–402, March 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-2-361.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-2-
                                        361.pdf.

Barney:1959:ETA

[BHD59] H. L. Barney, F. E. Haworth, and H. K. Dunn. An experimental transistorized artificial larynx. The Bell System Technical Journal, 38(6):1337–1356, November 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1337.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1337.pdf.

Buhrendorf:1956:LMM

[BHM56] F. G. Buhrendorf, H. A. Henning, and O. J. Murphy. A laboratory model magnetic drum translator for toll switching offices. The Bell System Technical Journal, 35(3):707–745, May 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-3-707.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-3-
707.pdf.

Busch:1955:MPM

[BHN55] K. J. Busch, A. D. Hasley, and Carl Neitzert. Magnetic pulse modulators. The Bell System Technical Journal, 34(5):943–993, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-5-943.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol34-1955/articles/bstj34-5-
                                        943.pdf.

Bouton:1955:EEA

[BHP55] G. M. Bouton, J. H. Heiss, and G. S. Phipps. Experimental extrusion of aluminum cable sheath at Bell Telephone Laboratories. The Bell System Technical Journal, 34(3):529–561, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-3-529.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-3-
                                        529.pdf.

Boyd:1957:NCS

[BHP57] R. C. Boyd, J. D. Howard, Jr., and L. Pedersen. A new carrier system for rural service. The Bell System Technical Journal, 36 (2):349–390, March 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-2-349.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-2-
                                        349.pdf.

Becker:1959:EVC

[BHW59] F. K. Becker, J. R. Hefele, and W. T. Wintringham. An experimental visual communication system. The Bell System Technical Journal, 38(1):141–176, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol38/
bstj38-1-141.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-1-
141.pdf.

Biggs:1951:DOP

[Big51] B. S. Biggs. Deterioration of organic polymers. The Bell System Technical Journal, 30(4): 1078–1102, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1078.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1078.pdf.

Blecher:1956:TCA

[Ble56] Franklin H. Blecher. Transistor circuits for analog and digital systems. The Bell System Technical Journal, 35(2):295–332, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-2-295.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-2-
295.pdf.

Blount:1953:TOU

[Blo53] F. E. Blount. Transistor oscillator for use in multifrequency pulsing current supply. The Bell System Technical Journal, 32(6): 1313–1331, November 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-6-1313.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-6-
                                        1313.pdf.

Bowen:1950:NMT

[BM50] A. E. Bowen and W. W. Mumford. A new microwave triode: Its performance as a modulator and as an amplifier. The Bell System Technical Journal, 29(4):531–552, October 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-4-531.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-4-
                                        531.pdf.

Bozorth:1951:FDE

[BMM51] R. M. Bozorth, W. P. Mason, and H. J. McSkimin. Frequency dependence of elastic constants and losses in nickel. The Bell System Technical Journal, 30(4):970–989, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-4-970.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-4-
                                        970.pdf.

Bobeck:1957:NSE

[Bob57] Andrew H. Bobeck. A new storage element suitable for large-sized memory arrays — the twistor. The Bell System Technical Journal, 36(6):1319–1340, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-6-1319.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-6-
1319.pdf.

Bozorth:1950:FAM

[Boz50] R. M. Bozorth. Factors affecting magnetic quality. The Bell System Technical Journal, 29 (2):251–286, April 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-2-251.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-2-
251.pdf.

Brown:1953:BPM

[BP53] J. T. L. Brown and C. E. Pollard.

Balanced polar mercury contact relay. The Bell System Technical Journal, 32(6):1393–1411, November 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-6-1393.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-6-
1393.pdf.

Babcock:1956:COW

[BRT56] W. C. Babcock, Esther Rentrop, and C. S. Thaeler. Crosstalk on open-wire lines. The Bell System Technical Journal, 35(2): 515–518, March 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-2-515.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-2-
                                        515.pdf.

Bennett:1956:SCE

[BS56a] D. C. Bennett and B. Sawyer. Single crystals of exceptional perfection and uniformity by zone leveling. The Bell System Technical Journal, 35(3):637–660, May 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-3-637.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-3-
                                        637.pdf.

Bond:1956:UIM

[BS56b] W. L. Bond and F. M. Smits. The use of an interference microscope for measurement of extremely thin surface layers. The Bell System Technical Journal, 35 (5):1209–1221, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-1209.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1209.pdf.

Brown:1954:ADE

[BT54] C. B. Brown and M. E. Terry. The application of designed experiments to the card translator. The Bell System Technical Journal, 33(2):369–398, March 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-2-369.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-2-
369.pdf.

Blackman:1958:MPSa

[BT58a] R. B. Blackman and J. W. Tukey.

The measurement of power spectra from the point of view of communications engineering — Part I. The Bell System Technical Journal, 37(1):185–282, January 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-1-185.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-1-
185.pdf.

Blackman:1958:MPSb

[BT58b] R. B. Blackman and J. W. Tukey.

The measurement of power spectra from the point of view of communications engineering — Part II. The Bell System Technical Journal, 37(2):485–569, March 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-2-485.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-2-
485.pdf.

Bullington:1953:FEM

[Bul53] Kenneth Bullington. Frequency economy in mobile radio bands. The Bell System Technical Journal, 32(1):42–62, January 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol32/bstj32-
                                        1-42.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol32-1953/
                                        articles/bstj32-1-42.pdf.

Bullington:1957:RPF

[Bul57] Kenneth Bullington. Radio propagation fundamentals. The Bell System Technical Journal, 36 (3):593–626, May 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-3-593.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-3-
                                        593.pdf.

Burford:1958:NVE

[Bur58] T. M. Burford. Nonstationary velocity estimation. The Bell System Technical Journal, 37 (4):1009–1021, July 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-4-1009.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-4-
                                        1009.pdf.

Burke:1959:EDD

[Bur59] Paul J. Burke. Equilibrium delay distribution for one channel with constant holding time, Poisson input and random service. The Bell System Technical Journal, 38(4):1021–1031, July 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-4-1021.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-4-
1021.pdf.

Brockbank:1957:RDN

[BWW57] R. A. Brockbank, D. C. Walker, and V. G. Welsby. Repeater design for the Newfoundland–Nova Scotia link. The Bell System Technical Journal, 36(1):245–276, January 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-1-245.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-1-
245.pdf.

Calbick:1951:IRE

[Cal51] C. J. Calbick. Inorganic replication in electron microscopy. The Bell System Technical Journal, 30 (4):798–824, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-798.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol30-1951/articles/bstj30-4-
                                        798.pdf.

Caruthers:1951:TCT

[Car51] R. S. Caruthers. The type N-1 carrier telephone system: Objectives and transmission features. The Bell System Technical Journal, 30(1):1–32, January 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol30-
                                        1951/articles/bstj30-1-1.pdf.

Crawford:1956:FDP

[CFJ56] A. B. Crawford, H. T. Friis, and W. C. Jakes, Jr. A 60-foot diameter parabolic antenna for propagation studies. The Bell System Technical Journal, 35(5):1199–1208, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-1199.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1199.pdf.

Crawford:1956:MAA

[CH56] A. B. Crawford and D. C. Hogg. Measurement of atmospheric attenuation at millimeter wavelengths. The Bell System Technical Journal, 35(4):907–916, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-4-907.pdf;   http://

www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-4-
907.pdf.

Crawford:1959:STP

[CHK59] A. B. Crawford, D. C. Hogg, and W. H. Kummer. Studies in tropospheric propagation beyond the horizon. The Bell System Technical Journal, 38(5):1067–1178, September 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-5-1067.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-5-
1067.pdf.

Chase:1954:TJD

[CHS54] F. H. Chase, B. H. Hamilton, and D. H. Smith. Transistors and junction diodes in telephone power plants. The Bell System Technical Journal, 33(4):827–858, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-827.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
827.pdf.

Crawford:1952:SFM

[CJ52] A. B. Crawford and W. C. Jakes, Jr. Selective fading of microwaves. The Bell System Technical Journal, 31(1):68–90, January 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol31/bstj31-

                                        1-68.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol31-1952/
                                        articles/bstj31-1-68.pdf.

Cook:1956:CH

[CKQ56] J. S. Cook, R. Kompfner, and C. F. Quate. Coupled helices. The Bell System Technical Journal, 35(1):127–178, January 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-1-127.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-1-
                                        127.pdf.

Clogston:1951:RSE

[Clo51] A. M. Clogston. Reduction of skin effect losses by the use of laminated conductors. The Bell System Technical Journal, 30(3):491–529, July 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-3-491.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-3-
                                        491.pdf.

Clos:1953:SNB

[Clo53] Charles Clos. A study of nonblocking switching networks. The Bell System Technical Journal, 32 (2):406–424, March 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-2-406.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol32-1953/articles/bstj32-2-
406.pdf.

Clogston:1955:RPF

[Clo55] A. M. Clogston. Relaxation phenomena in ferrites. The Bell System Technical Journal, 34 (4):739–760, July 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-4-739.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-4-
739.pdf.

Clark:1952:ASN

[CO52] A. B. Clark and H. S. Osborne.

Automatic switching for nationwide telephone service. The Bell System Technical Journal, 31(5): 823–831, September 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-823.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
823.pdf.

Colley:1953:EWPa

[Col53a] Reginald H. Colley. The evaluation of wood preservatives: Part I — interpretation and correlation of the results of laboratory soil-block tests and outdoor test plot experience, with special reference to oil-type materials. The Bell System Technical Journal, 32(1):120–169, January 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-1-120.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-1-
                                        120.pdf.

Colley:1953:EWPb

[Col53b] Reginald H. Colley. The evaluation of wood preservatives: Part II. The Bell System Technical Journal, 32(2):425–505, March 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-2-425.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-2-
                                        425.pdf.

Cook:1955:TVC

[Coo55] J. S. Cook. Tapered velocity couplers. The Bell System Technical Journal, 34(4):807–822, July 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-4-807.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-4-
                                        807.pdf.

Curtis:1957:IID

[CR57] H. E. Curtis and S. O. Rice. Interchannel interference due to klystron pulling. The Bell System Technical Journal, 36(3):645–652, May 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol36/
bstj36-3-645.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-3-
645.pdf.

Coolidge:1959:ART

[CR59] O. H. Coolidge and G. C. Reier.

An appraisal of received telephone speech volume. The Bell System Technical Journal, 38 (3):877–897, May 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-3-877.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-3-
877.pdf.

Cole:1956:AMT

[CS56] C. C. Cole and H. R. Shillington. Automatic machine for testing capacitors and resistancecapacitance networks. The Bell System Technical Journal, 35(5): 1179–1198, September 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-5-1179.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-5-
1179.pdf.

Cutler:1956:NPS

[Cut56] C. C. Cutler. The nature of power saturation in traveling wave tubes. The Bell System Technical Journal, 35(4):841–876, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-4-841.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-4-
                                        841.pdf.

Clos:1952:DHT

[CW52] Charles Clos and Roger I. Wilkinson. Dialing habits of telephone customers. The Bell System Technical Journal, 31(1): 32–67, January 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol31/bstj31-
                                        1-32.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol31-1952/
                                        articles/bstj31-1-32.pdf.

Darlington:1950:RCP

[Dar50] Sidney Darlington. Realization of a constant phase difference. The Bell System Technical Journal, 29 (1):94–104, January 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol29/bstj29-
                                        1-94.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol29-1950/
                                        articles/bstj29-1-94.pdf.

Darlington:1951:PAM

[Dar51a] Sidney Darlington. The potential analogue method of network synthesis. The Bell System Technical Journal, 30(2):315–365, April 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol30/
bstj30-2-315.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-2-
315.pdf.

Darrow:1951:SWC

[Dar51b] Karl K. Darrow. The scientific work of C. J. Davisson. The Bell System Technical Journal, 30 (4):786–797, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-786.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
786.pdf.

Darlington:1952:NSU

[Dar52] Sidney Darlington. Network synthesis using Tchebycheff polynomial series. The Bell System Technical Journal, 31(4):613–665, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-4-613.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
613.pdf.

Darrow:1953:MRPa

[Dar53a] Karl K. Darrow. Magnetic resonance: Part I — nuclear magnetic resonance. The Bell System Technical Journal, 32(1): 74–99, January 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.

                                        com/BSTJ/images/Vol32/bstj32-
                                        1-74.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol32-1953/
                                        articles/bstj32-1-74.pdf.

Darrow:1953:MRPb

[Dar53b] Karl K. Darrow. Magnetic resonance: Part II — magnetic resonance of electrons. The Bell System Technical Journal, 32(2): 384–405, March 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-2-384.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-2-
                                        384.pdf.

Darlington:1958:LLS

[Dar58] Sidney Darlington. Linear leastsquares smoothing and prediction, with applications. The Bell System Technical Journal, 37 (5):1221–1294, September 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-5-1221.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-5-
                                        1221.pdf.

Dawson:1957:EDP

[Daw57] R. W. Dawson. An experimental dual polarization antenna feed for three radio relay bands. The Bell System Technical Journal, 36(2):391–408, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-2-391.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-2-
391.pdf.

Davis:1954:ERT

[DB54] T. E. Davis and A. L. Blaha.

Electronic relay transfer. The Bell System Technical Journal, 33(4):925–938, July 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-925.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
925.pdf.

DeLange:1952:PSM

[De 52] O. E. De Lange. Propagation studies at microwave frequencies by means of very short pulses. The Bell System Technical Journal, 31 (1):91–103, January 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol31/bstj31-
1-91.pdf;   http://www.alcatel-
lucent.com/bstj/vol31-1952/
articles/bstj31-1-91.pdf.

DeLange:1958:THS

[De 58] O. E. De Lange. The timing of high-speed regenerative repeaters. The Bell System Technical Journal, 37(6):1455–1486, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-6-1455.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-6-
                                        1455.pdf.

DeLange:1956:ERB

[DeL56] O. E. DeLange. Experiments on the regeneration of binary microwave pulses. The Bell System Technical Journal, 35(1): 67–90, January 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol35/bstj35-
                                        1-67.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol35-1956/
                                        articles/bstj35-1-67.pdf.

Desoer:1957:NCP

[Des57] C. A. Desoer. A network containing a periodically operated switch solved by successive approximations. The Bell System Technical Journal, 36(6):1403–1428, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-6-1403.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-6-
                                        1403.pdf.

Doba:1955:NLV

[DK55] Stephen Doba, Jr. and A. Robert Kolding. A new local video transmission system. The Bell System Technical Journal, 34 (4):677–712, July 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-4-677.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-4-
677.pdf.

Dodge:1953:LCS

[DKK53] H. F. Dodge, B. J. Kinsburg, and M. K. Kruger. The L3 coaxial system: Quality control requirements. The Bell System Technical Journal, 32(4):943–967, July 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-4-943.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-4-
943.pdf.

DeLange:1958:ETR

[DP58] O. E. De Lange and M. Pustelnyk. Experiments on the timing of regenerative repeaters. The Bell System Technical Journal, 37(6):1487–1500, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-6-1487.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-6-
1487.pdf.

Dacey:1955:FET

[DR55] G. C. Dacey and I. M. Ross.

The field effect transistor. The Bell System Technical Journal, 34(6):1149–1189, November 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-6-1149.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-6-
                                        1149.pdf.

Danielson:1956:DAB

[DRS56] W. E. Danielson, J. L. Rosenfeld, and J. A. Saloom. A detailed analysis of beam formation with electron guns of the Pierce type. The Bell System Technical Journal, 35(2):375–420, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-2-375.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-2-
                                        375.pdf.

DeGrasse:1959:TLS

[DSDS59] R. W. DeGrasse, E. O. Schulz-DuBois, and H. E. D. Scovil. The three-level solid state travelingwave maser. The Bell System Technical Journal, 38(2): 305–334, March 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-2-305.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-2-
                                        305.pdf.

Early:1953:DTJ

[Ear53] J. M. Early. Design theory of junction transistors. The Bell System Technical Journal, 32(6): 1271–1312, November 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-6-1271.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-6-
1271.pdf.

Early:1954:PPP

[Ear54] J. M. Early. P-N-I-P and N-P-I-N junction transistor triodes. The Bell System Technical Journal, 33(3):517–533, May 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-3-517.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-3-
517.pdf.

Earle:1955:NMT

[Ear55] Norman E. Earle. New manufacturing techniques for precision transformers for the L3 coaxial carrier system. The Bell System Technical Journal, 34(2):291–307, March 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-2-291.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-2-
291.pdf.

Edwards:1956:FCM

[Edw56] C. F. Edwards. Frequency conversion by means of a nonlinear admittance. The Bell System Technical Journal, 35(6):1403–1416, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-6-1403.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-6-
1403.pdf.

Elmendorf:1953:LCS

[EEKG53] C. H. Elmendorf, R. D. Ehrbar, R. H. Klie, and A. J. Grossman. The L3 coaxial system: System design. The Bell System Technical Journal, 32(4):781–832, July 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-4-781.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-4-
781.pdf.

Eggleston:1950:CCL

[Egg50] Richard C. Eggleston. Conversion of concentrated loads on wood crossarms to loads distributed at each pin position. The Bell System Technical Journal, 29(1): 105–116, January 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-1-105.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-1-
105.pdf.

Elmendorf:1957:OIE

[EH57] C. H. Elmendorf and Bruce C.

Heezen. Oceanographic information for engineering submarine cable systems. The Bell System Technical Journal, 36(5): 1047–1093, September 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-5-1047.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-5-
1047.pdf.

Ellwood:1955:MTP

[Ell55] W. B. Ellwood. The measurement of the transient power and energy dissipated in closing switch contacts. The Bell System Technical Journal, 34(6):1191–1223, November 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-6-1191.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-6-
1191.pdf.

Elliott:1959:ESW

[Ell59] S. J. Elliott. Evaluation of solderless wrapped connections for central office use. The Bell System Technical Journal, 38(4): 1033–1059, July 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-4-1033.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol38-1959/articles/bstj38-4-
                                        1033.pdf.

Edwards:1952:TCS

[EM52] Paul G. Edwards and L. R. Montfort. The type-O carrier system. The Bell System Technical Journal, 31(4):688–723, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-4-688.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-4-
                                        688.pdf.

Ebers:1955:DAJ

[EM55] J. J. Ebers and S. L. Miller. Design of alloyed junction germanium transistors for high-speed switching. The Bell System Technical Journal, 34(4):761–781, July 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-4-761.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-4-
                                        761.pdf.

Eppler:1954:TMC

[Epp54] W. T. Eppler. Thickness measurement and control in the manufacture of polyethylene cable sheath. The Bell System Technical Journal, 33(3):559–577, May 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-3-559.pdf;   http://

www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-3-
559.pdf.

Evans:1958:OTA

[Eva58] David H. Evans. Optimum tolerance assignment to yield minimum manufacturing cost. The Bell System Technical Journal, 37 (2):461–484, March 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-2-461.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-2-
461.pdf.

Friis:1957:RTP

[FCH57] H. T. Friis, A. B. Crawford, and D. C. Hogg. A reflection theory for propagation beyond the horizon. The Bell System Technical Journal, 36(3):627–644, May 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-3-627.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-3-
627.pdf.

Feher:1957:SCM

[Feh57] G. Feher. Sensitivity considerations in microwave paramagnetic resonance absorption techniques. The Bell System Technical Journal, 36(2):449–484, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-2-449.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-2-
                                        449.pdf.

Feiner:1958:HSL

[FG58] A. Feiner and L. F. Goeller, Jr. A high-speed line scanner for use in an electronic switching system. The Bell System Technical Journal, 37(6):1383–1403, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-6-1383.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-6-
                                        1383.pdf.

Frost:1953:TMT

[FKR53] G. R. Frost, William Keister, and Alistair E. Ritchie. A throwdown machine for telephone traffic studies. The Bell System Technical Journal, 32(2):292–359, March 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-2-292.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-2-
                                        292.pdf.

Felder:1956:ITN

[FL56] H. H. Felder and E. N. Little. Intertoll trunk net loss maintenance under operator distance and direct distance dialing. The Bell System Technical Journal, 35(4):955–972, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-4-955.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-4-
955.pdf.

Forster:1956:EST

[FM56] J. H. Forster and L. E. Miller.

The effect of surface treatments on point-contact transistor characteristics. The Bell System Technical Journal, 35(4):767–811, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-4-767.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-4-

767.pdf. See erratum [Ano57m].

Fox:1955:BAF

[FMW55] A. G. Fox, S. E. Miller, and M. T.

Weiss. Behavior and applications of ferrites in the microwave region. The Bell System Technical Journal, 34(1):5–103, January 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-1-5.pdf;   http://www.
alcatel-lucent.com/bstj/vol34-
1955/articles/bstj34-1-5.pdf.

Fox:1955:WCW

[Fox55] A. G. Fox. Wave coupling by warped normal modes. The Bell System Technical Journal, 34(4):823–852, July 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-4-823.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-4-
                                        823.pdf.

Felder:1956:ATT

[FPS56] H. H. Felder, A. J. Pascarella, and H. F. Shoffstall. Automatic testing of transmission and operational functions of intertoll trunks. The Bell System Technical Journal, 35(4):927–954, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-4-927.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-4-
                                        927.pdf.

Feldman:1958:ABD

[FR58] T. Feldman and J. W. Rieke. Application of breakdown devices to large multistage switching networks. The Bell System Technical Journal, 37(6):1421–1453, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-6-1421.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-6-
                                        1421.pdf.

Freimanis:1958:SDE

[Fre58] L. Freimanis. A signal distributor for electronic switching systems. The Bell System Technical Journal, 37(6):1405–1419, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-6-1405.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-6-
1405.pdf.

Fry:1952:IDF

[Fry52] J. R. Fry. Important design factors influencing reliability of relays. The Bell System Technical Journal, 31(5):976–998, September 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-976.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
976.pdf.

Fuchs:1959:WBD

[Fuc59] F. J. Fuchs, Jr. Waveguide bending design analysis: Theory of bending and formulae for determination of wall thicknesses. The Bell System Technical Journal, 38(6):1457–1484, November 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1457.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1457.pdf.

Ford:1951:DET

[FW51] G. T. Ford and E. J. Walsh. The development of electron tubes for a new coaxial transmission system. The Bell System Technical Journal, 30(4):1103–1128, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-4-1103.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-4-
                                        1103.pdf.

Galt:1954:MID

[Gal54] J. K. Galt. Motion of individual domain walls in a nickel–iron ferrite. The Bell System Technical Journal, 33(5):1023–1054, September 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-5-1023.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-5-
                                        1023.pdf.

Galt:1955:MID

[Gal55] J. K. Galt. Motion of individual domain walls in a nickel–iron ferrite. The Bell System Technical Journal, 34(2):439–441, March 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-2-439.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-2-
                                        439.pdf.

Gallaher:1959:BPS

[Gal59] L. E. Gallaher. Beam-positioning servo system for the flying spot store. The Bell System Technical Journal, 38(2):425–444, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-425.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
425.pdf.

Garrett:1956:DCS

[GB56] C. G. B. Garrett and W. H.

Brattain. Distribution and crosssections of fast states on germanium surfaces. The Bell System Technical Journal, 35(5):1041– 1058, September 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-5-1041.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-5-
1041.pdf.

Gianola:1959:LMD

[GC59] U. F. Gianola and T. H. Crowley. The laddic — A magnetic device for performing logic. The Bell System Technical Journal, 38 (1):45–72, January 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol38/bstj38-
1-45.pdf;   http://www.alcatel-
lucent.com/bstj/vol38-1959/
articles/bstj38-1-45.pdf.

Germer:1951:HDE

[Ger51] L. H. Germer. Heat dissipation at the electrodes of a short electric arc. The Bell System Technical Journal, 30(4):933–944, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-4-933.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-4-
                                        933.pdf.

Gilmore:1958:NDM

[GGI58] A. C. Gilmore, P. R. Gray, and W. S. Irvine. New developments in military switching. The Bell System Technical Journal, 37(2):375–400, March 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-2-375.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-2-
                                        375.pdf.

Gilbert:1951:TSC

[Gil51a] E. N. Gilbert. N-terminal switching circuits. The Bell System Technical Journal, 30(3):668–688, July 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-3-668.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-3-
                                        668.pdf.

Gilbert:1951:STC

[Gil51b] J. J. Gilbert. A submarine telephone cable with submerged repeaters. The Bell System Technical Journal, 30 (1):65–87, January 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol30/bstj30-
1-65.pdf;   http://www.alcatel-
lucent.com/bstj/vol30-1951/
articles/bstj30-1-65.pdf.

Gilbert:1952:CSA

[Gil52] E. N. Gilbert. A comparison of signalling alphabets. The Bell System Technical Journal, 31(3):504–522, May 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-3-504.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-3-
504.pdf.

Gilbert:1958:GCP

[Gil58] E. N. Gilbert. Gray codes and paths on the n-cube. The Bell System Technical Journal, 37(3):815–826, May 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-815.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-3-
815.pdf.

Gleichmann:1957:RDN

[GLWB57] T. F. Gleichmann, A. H. Lince, M. C. Wooley, and F. J. Braga. Repeater design for the North Atlantic Link. The Bell System Technical Journal, 36(1): 69–101, January 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol36/bstj36-
1-69.pdf;   http://www.alcatel-
lucent.com/bstj/vol36-1957/
articles/bstj36-1-69.pdf.

Gray:1953:DME

[GM53] A. N. Gray and G. E. Murray. Development and manufacture of electroformed conductor for telephone drop wire. The Bell System Technical Journal, 32 (5):1099–1135, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-5-1099.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-5-
1099.pdf.

Gammie:1955:SNI

[GM55a] J. Gammie and J. L. Merrill, Jr. Stability of negative impedance elements in short transmission lines. The Bell System Technical Journal, 34(2):333–360, March 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-2-333.pdf;   http://
www.alcatel-lucent.com/bstj/

vol34-1955/articles/bstj34-2-
333.pdf.

Gilbert:1955:ODD

[GM55b] E. N. Gilbert and S. P. Morgan.

Optimum design of directive antenna arrays subject to random variations. The Bell System Technical Journal, 34(3):637–663, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-3-637.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-3-
637.pdf.

Gilbert:1959:VLB

[GM59] E. N. Gilbert and E. F. Moore.

Variable-length binary encodings. The Bell System Technical Journal, 38(4):933–967, July 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-4-933.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-4-
933.pdf.

Goodall:1951:TPC

[Goo51] W. M. Goodall. Television by pulse code modulation. The Bell System Technical Journal, 30 (1):33–49, January 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol30/bstj30-
1-33.pdf;   http://www.alcatel-

lucent.com/bstj/vol30-1951/
articles/bstj30-1-33.pdf.

Gilbert:1957:CPP

[GP57] E. N. Gilbert and H. O. Pollak. Coincidences in Poisson patterns. The Bell System Technical Journal, 36(4):1005–1033, July 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-4-1005.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-4-
1005.pdf.

Geusic:1959:PRS

[GPSD59] J. E. Geusic, Martin Peter, and E. O. Schulz-DuBois. Paramagnetic resonance spectrum of Cr+++ in emerald. The Bell System Technical Journal, 38(1):291–296, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-1-291.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-1-
291.pdf.

Grisdale:1951:PFR

[GPV51] R. O. Grisdale, A. C. Pfister, and W. Van Roosbroeck. Pyrolytic film resistors: Carbon and borocarbon. The Bell System Technical Journal, 30(2):271–314, April 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-2-271.pdf;   http://

www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-2-
271.pdf.

Gray:1951:ESD

[Gra51] Frank Gray. Electron streams in a diode. The Bell System Technical Journal, 30(4):830–854, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-830.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
830.pdf.

Gramels:1953:SRF

[Gra53] J. Gramels. Selenium rectifiers– factors in their application. The Bell System Technical Journal, 32(6):1469–1492, November 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-6-1469.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-6-
1469.pdf.

Green:1958:T

[Gre58] E. I. Green. Telephone. The Bell System Technical Journal, 37 (2):289–338, March 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-2-289.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-2-
289.pdf.

Grubbs:1959:HED

[Gru59] W. J. Grubbs. Hall effect devices. The Bell System Technical Journal, 38(3):853–876, May 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-3-853.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-3-
                                        853.pdf.

Germer:1957:AEC

[GS57] L. H. Germer and J. L. Smith. Activation of electrical contacts by organic vapors. The Bell System Technical Journal, 36 (3):769–812, May 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-3-769.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-3-
                                        769.pdf.

Greenwood:1958:HSB

[GS58] T. S. Greenwood and R. E. Staehler. A high-speed barrier grid store. The Bell System Technical Journal, 37(5):1195–1220, September 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-5-1195.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-5-
                                        1195.pdf.

Garrett:1953:LCS

[GTW53] R. F. Garrett, T. L. Tuffnell, and R. A. Waddell. The L3 coaxial system: Application of quality control requirements in the manufacture of components. The Bell System Technical Journal, 32(4):969–1005, July 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-4-969.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-4-
969.pdf.

Hagelbarger:1959:RCE

[Hag59] D. W. Hagelbarger. Recurrent codes: easily mechanized, burstcorrecting, binary codes. The Bell System Technical Journal, 38(4):969–984, July 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-4-969.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-4-
969.pdf.

Hamming:1950:EDE

[Ham50] R. W. Hamming. Error detecting and error correcting codes. The Bell System Technical Journal, 29(2):147–160, April 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-2-147.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol29-1950/articles/bstj29-2-
                                        147.pdf.

Hansen:1956:AMT

[Han56] L. D. Hansen. Automatic manufacturing testing of relay switching circuits. The Bell System Technical Journal, 35(5):1155–1178, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-1155.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1155.pdf.

Hartley:1950:MMM

[Har50a] R. V. L. Hartley. Matter, A mode of motion. The Bell System Technical Journal, 29(3):350–368, July 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-3-350.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-3-
                                        350.pdf.

Hartley:1950:RDW

[Har50b] R. V. L. Hartley. The reflection of diverging waves by a gyrostatic medium. The Bell System Technical Journal, 29(3):369–389, July 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-3-369.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-3-
                                        369.pdf.

Harrison:1952:ELP

[Har52] C. W. Harrison. Experiments with linear prediction in television. The Bell System Technical Journal, 31(4):764–783, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-4-764.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
764.pdf.

Harkless:1959:NCR

[Har59] Earl T. Harkless. A network for combining radio systems at 4, 6 and 11 kmc. The Bell System Technical Journal, 38(5):1253–1267, September 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-5-1253.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-5-
1253.pdf.

Hayward:1952:RTT

[Hay52] W. S. Hayward, Jr. The reliability of telephone traffic load measurements by switch counts. The Bell System Technical Journal, 31(2):357–377, March 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-2-357.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-2-
357.pdf.

Halsey:1957:SDN

[HB57] R. J. Halsey and J. F. Bampton. System design for the Newfoundland–Nova Scotia link. The Bell System Technical Journal, 36(1):217–244, January 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-1-217.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-1-
217.pdf.

Hines:1955:DMB

[HCM55] M. E. Hines, M. Chruney, and J. A. McCarthy. Digital memory in barrier-grid storage tubes. The Bell System Technical Journal, 34(6):1241–1264, November 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-6-1241.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-6-
1241.pdf.

Hermance:1958:ODP

[HE58] H. W. Hermance and T. F. Egan. Organic deposits on precious metal contacts. The Bell System Technical Journal, 37(3):739–776, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-739.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-3-
739.pdf.

Hathaway:1959:RAK

[HE59] S. D. Hathaway and H. W. Evans.

Radio attenuation at 11 kmc and some implications affecting relay system engineering. The Bell System Technical Journal, 38 (1):73–97, January 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol38/bstj38-
1-73.pdf;   http://www.alcatel-
lucent.com/bstj/vol38-1959/
articles/bstj38-1-73.pdf.

Heidenreich:1951:ETT

[Hei51] R. D. Heidenreich. Electron transmission through thin metal sections with application to selfrecovery in cold worked aluminum. The Bell System Technical Journal, 30(4):867–887, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-867.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
867.pdf.

Helm:1959:T

[Hel59] H. A. Helm. The Z transformation. The Bell System Technical Journal, 38(1):177–196, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-1-177.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-1-
177.pdf.

Herring:1955:TPM

[Her55] Conyers Herring. Transport properties of a many-valley semiconductor. The Bell System Technical Journal, 34(2):237–290, March 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-2-237.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-2-
                                        237.pdf.

Herring:1959:APD

[HGK59] C. Herring, T. H. Geballe, and J. E. Kunzler. Analysis of phonon-drag thermomagnetic effects in n-type germanium. The Bell System Technical Journal, 38(3):657–747, May 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-3-657.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-3-
                                        657.pdf.

Hoover:1959:SDF

[HHH59] C. W. Hoover, Jr., G. Haugk, and D. R. Herriott. System design of the flying spot store. The Bell System Technical Journal, 38(2):365–401, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-2-365.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-2-
                                        365.pdf.

Hines:1950:WRM

[Hin50] M. E. Hines. A wide range microwave sweeping oscillator. The Bell System Technical Journal, 29 (4):553–559, October 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-4-553.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-4-
553.pdf.

Hampton:1953:CTN

[HN53] L. N. Hampton and J. B. Newsom.

The card translator for nationwide dialing. The Bell System Technical Journal, 32(5):1037–1098, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-5-1037.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-5-
1037.pdf.

Hogan:1952:FFE

[Hog52] C. L. Hogan. The ferromagnetic Faraday Effect at microwave frequencies and its applications: The microwave gyrator. The Bell System Technical Journal, 31 (1):1–31, January 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol31-
1952/articles/bstj31-1-1.pdf.

Holden:1951:LME

[Hol51] A. N. Holden. Longitudinal modes of elastic waves in isotropic cylinders and slabs. The Bell System Technical Journal, 30(4): 956–969, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-956.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
956.pdf.

Hopkins:1952:PSC

[Hop52] I. L. Hopkins. Principle strains in cable sheaths and other buckled surfaces. The Bell System Technical Journal, 31(3):523–529, May 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-3-523.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-3-
523.pdf.

Hovgaard:1955:DRS

[HP55] O. M. Hovgaard and G. E. Perreault. Development of reed switches and relays. The Bell System Technical Journal, 34(2): 309–332, March 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-2-309.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-2-
309.pdf.

Hysko:1952:CTS

[HRR52] J. L. Hysko, W. T. Rea, and L. C. Roberts. A carrier telegraph system for short-haul applications. The Bell System Technical Journal, 31(4):666–687, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-4-666.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
666.pdf.

Hohn:1955:BMD

[HS55] Franz E. Hohn and L. Robert Schissler. Boolean matrices and the design of combinational relay switching circuits. The Bell System Technical Journal, 34(1): 177–202, January 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-1-177.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-1-
177.pdf.

Hoover:1958:FCD

[HSK58] C. W. Hoover, Jr., R. E. Staehler, and R. W. Ketchledge. Fundamental concepts in the design of the flying spot store. The Bell System Technical Journal, 37 (5):1161–1194, September 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-5-1161.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-5-
                                        1161.pdf.

Huntley:1953:TDI

[Hun53] H. R. Huntley. Transmission design of intertoll telephone trunks. The Bell System Technical Journal, 32(5):1019–1036, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-5-1019.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-5-
                                        1019.pdf.

Hussey:1953:SDG

[Hus53] Luther W. Hussey. Semiconductor diode gates. The Bell System Technical Journal, 32(5): 1137–1154, September 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-5-1137.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-5-
                                        1137.pdf.

Horton:1955:TDI

[HV55] A. W. Horton, Jr. and H. E. Vaughan. Transmission of digital information over telephone circuits. The Bell System Technical Journal, 34(3):511–528, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-3-511.pdf;   http://

www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-3-
511.pdf.

Inglis:1951:ITS

[IT51] A. H. Inglis and W. L. Tuffnell.

An improved telephone set. The Bell System Technical Journal, 30 (2):239–270, April 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-2-239.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-2-
239.pdf.

Jacobs:1958:NTL

[Jac58] Ira Jacobs. The nonuniform transmission line as a broadband termination. The Bell System Technical Journal, 37(4):913–924, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-4-913.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-4-
913.pdf.

Jensen:1951:DCR

[Jen51] A. G. Jensen. The Davisson cathode ray television tube using deflection modulation. The Bell System Technical Journal, 30 (4):855–866, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-855.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-4-
                                        855.pdf.

Jack:1957:RSC

[JLL57] J. S. Jack, Capt. W. H. Leech, and H. A. Lewis. Route selection and cable laying for the transatlantic cable system. The Bell System Technical Journal, 36 (1):293–326, January 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-1-293.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-1-
                                        293.pdf.

Joel:1956:ETS

[Joe56a] A. E. Joel. Electronics in telephone switching systems. The Bell System Technical Journal, 35(5):991–1018, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-991.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        991.pdf.

Joel:1956:ERC

[Joe56b] A. E. Joel, Jr. An experimental remote controlled line concentrator. The Bell System Technical Journal, 35(2):249–293, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/

bstj35-2-249.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-2-
249.pdf.

Joel:1958:ESS

[Joe58] A. E. Joel, Jr. An experimental switching system using new electronic techniques. The Bell System Technical Journal, 37(5): 1091–1124, September 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-5-1091.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-5-
1091.pdf.

Johnston:1954:ITC

[Joh54] D. F. Johnston. Intertoll trunk concentrating equipment. The Bell System Technical Journal, 33 (2):303–327, March 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-2-303.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-2-
303.pdf.

Julesz:1959:MCT

[Jul59] Bela Julesz. A method of coding television signals based on edge detection. The Bell System Technical Journal, 38(4):1001–1020, July 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/

                                        bstj38-4-1001.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-4-
                                        1001.pdf.

Kaplan:1955:SDS

[Kap55] E. L. Kaplan. Signal-detection studies, with applications. The Bell System Technical Journal, 34 (2):403–437, March 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-2-403.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-2-
                                        403.pdf.

Kaylor:1953:SSS

[Kay53] R. L. Kaylor. A statistical study of selective fading of super-high frequency radio signals. The Bell System Technical Journal, 32 (5):1187–1202, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-5-1187.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-5-
                                        1187.pdf.

Kelly:1951:DCJ

[Kel51] M. J. Kelly. Dr. C. J. Davisson. The Bell System Technical Journal, 30(4):779–785, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-4-779.pdf;   http://

www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
779.pdf.

Keller:1952:NGP

[Kel52] Arthur C. Keller. A new general purpose relay for telephone switching systems. The Bell System Technical Journal, 31(6): 1023–1067, November 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-6-1023.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-6-
1023.pdf.

Kelly:1956:NII

[Kel56] J. L. Kelly, Jr. A new interpretation of information rate. The Bell System Technical Journal, 35(4):917–926, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-4-917.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-4-
917.pdf.

Kempf:1951:CIS

[Kem51] R. A. Kempf. Coaxial impedance standards. The Bell System Technical Journal, 30(3):689–705, July 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-3-689.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol30-1951/articles/bstj30-3-
                                        689.pdf.

Ketchledge:1955:DFA

[Ket55] R. W. Ketchledge. Distortion in feedback amplifiers. The Bell System Technical Journal, 34(6): 1265–1285, November 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-6-1265.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-6-
                                        1265.pdf.

Ketchledge:1959:LDS

[Ket59] R. W. Ketchledge. Logic for a digital servo system. The Bell System Technical Journal, 38 (1):1–17, January 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol38-
                                        1959/articles/bstj38-1-1.pdf.

Ketchledge:1953:LCS

[KF53] R. W. Ketchledge and T. R. Finch. The L3 coaxial system: Equalization and regulation. The Bell System Technical Journal, 32(4):833–878, July 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-4-833.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol32-1953/articles/bstj32-4-
833.pdf.

Keefer:1958:RCB

[KG58] H. J. Keefer and R. H. Gumley. Relay contact behavior under non-eroding circuit conditions. The Bell System Technical Journal, 37(3):777–814, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-777.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-3-
777.pdf.

Kock:1951:PMD

[KH51] W. E. Kock and F. K. Harvey. A photographic method for displaying sound wave and microwave space patterns. The Bell System Technical Journal, 30(3):564–587, July 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-3-564.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-3-
564.pdf.

Kohman:1955:SME

[KHD55] G. T. Kohman, H. W. Hermance, and G. H. Downes. Silver migration in electrical insulation. The Bell System Technical Journal, 34(6):1115–1147, November 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/

                                        bstj34-6-1115.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-6-
                                        1115.pdf.

King:1954:CAM

[Kin54] G. V. King. Centralized automatic message accounting system. The Bell System Technical Journal, 33(6):1331–1342, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-6-1331.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-6-
                                        1331.pdf.

King:1956:OMA

[Kin56] A. P. King. Observed 5-6 mm attenuation for the circular electric wave in small and medium-sized pipes. The Bell System Technical Journal, 35(5):1115–1128, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-1115.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1115.pdf.

Kinariwala:1959:SAR

[Kin59a] B. K. Kinariwala. Synthesis of active RC networks. The Bell System Technical Journal, 38(5): 1269–1316, September 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/

bstj38-5-1269.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-5-
1269.pdf.

King:1959:ANS

[Kin59b] J. C. King. The anelasticity of natural and synthetic quartz at low temperatures. The Bell System Technical Journal, 38(2): 573–602, March 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-573.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
573.pdf.

King:1959:ECT

[Kin59c] Mrs. R. A. King. An experimental clogston 2 transmission line. The Bell System Technical Journal, 38(2):517–536, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-517.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
517.pdf.

Kisliuk:1958:UCR

[Kis58] P. Kisliuk. Using contact resistance to measure adsorption of gases on metals. The Bell System Technical Journal, 37(4):925–949, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/

                                        bstj37-4-925.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-4-
                                        925.pdf.

Kleinman:1956:FCP

[Kle56] D. A. Kleinman. The forward characteristic of the PIN diode. The Bell System Technical Journal, 35(3):685–706, May 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-3-685.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-3-
                                        685.pdf.

King:1956:TLD

[KM56] A. P. King and E. A. Marcatili. Transmission loss due to resonance of loosely-coupled modes in a multi-mode system. The Bell System Technical Journal, 35(4):899–906, July 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-4-899.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-4-
                                        899.pdf.

Kingsbury:1952:PPI

[KO52] Edwin F. Kingsbury and Russell S. Ohl. Photoelectric properties of ionically bombarded silicon. The Bell System Technical Journal, 31(4):802–815, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-4-802.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
802.pdf.

Kahl:1951:SDF

[KP51] W. E. Kahl and L. Pedersen.

Some design features of the N-1 carrier telephone system. The Bell System Technical Journal, 30 (2):418–446, April 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-2-418.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-2-
418.pdf.

Kelly:1957:TCH

[KR57] Mervin J. Kelly and Sir Gordon Radley. Transatlantic communications — an historical resume. The Bell System Technical Journal, 36(1):1–5, January 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol36-
1957/articles/bstj36-1-1.pdf.

Kretzmer:1952:STS

[Kre52] E. R. Kretzmer. Statistics of television signals. The Bell System Technical Journal, 31(4):751–763, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-4-751.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-4-
                                        751.pdf.

Kronacher:1957:DPA

[Kro57] G. Kronacher. Design, performance and application of the vernier resolver. The Bell System Technical Journal, 36(6):1487–1500, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-6-1487.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-6-
                                        1487.pdf.

Lechleider:1959:MCJ

[Lec59] J. W. Lechleider. Mode conversion at the junction of helix waveguide and copper pipe. The Bell System Technical Journal, 38(5):1317–1329, September 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-5-1317.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-5-
                                        1317.pdf.

Lee:1955:ASN

[Lee55] C. Y. Lee. Analysis of switching networks. The Bell System Technical Journal, 34(6):1287– 1315, November 1955. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol34/
bstj34-6-1287.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-6-
1287.pdf.

Lee:1956:HFD

[Lee56] Charles A. Lee. A high-frequency diffused base germanium transistor. The Bell System Technical Journal, 35(1):23–34, January 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol35/bstj35-
1-23.pdf;   http://www.alcatel-
lucent.com/bstj/vol35-1956/
articles/bstj35-1-23.pdf.

Leenov:1958:GNF

[Lee58] D. Leenov. Gain and noise figure of a variable-capacitance upconverter. The Bell System Technical Journal, 37(4):989–1008, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-4-989.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-4-
989.pdf.

Lee:1959:RSC

[Lee59] C. Y. Lee. Representation of switching circuits by binarydecision programs. The Bell System Technical Journal, 38 (4):985–999, July 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/

                                        bstj38-4-985.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-4-
                                        985.pdf.

Levenbach:1958:CSD

[Lev58] G. J. Levenbach. Contribution of statistics to the development program of a transformer for L3 carrier system. The Bell System Technical Journal, 37 (1):23–54, January 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol37/bstj37-
                                        1-23.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol37-1958/
                                        articles/bstj37-1-23.pdf.

Lebert:1957:CDM

[LFB57] A. W. Lebert, H. B. Fischer, and M. C. Biskeborn. Cable design and manufacture for the transatlantic submarine cable system. The Bell System Technical Journal, 36(1):189–216, January 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-1-189.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-1-

189.pdf. See erratum [Ano57m].

Lamb:1957:RPN

[LH57] H. A. Lamb and W. W. Heffner. Repeater production for the North Atlantic Link. The Bell System Technical Journal, 36(1): 103–138, January 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-1-103.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-1-
103.pdf.

Leagus:1959:VLS

[LLM59] Dolores C. Leagus, C. Y. Lee, and George H. Mealy. Verification of the logic structure of an experimental switching system on a digital computer. The Bell System Technical Journal, 38(2):467–476, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-467.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
467.pdf.

Lloyd:1957:BBC

[Llo57] S. P. Lloyd. Binary block coding. The Bell System Technical Journal, 36(2):517–535, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-2-517.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-2-
517.pdf.

Lovell:1955:EPS

[LMM55] C. A. Lovell, J. H. McGuigan, and O. J. Murphy. An experimental polytonic signaling system. The Bell System Technical Journal, 34(4):783–806, July 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-4-783.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-4-
                                        783.pdf.

Laico:1956:MPT

[LMM56] J. P. Laico, H. L. McDowell, and C. R. Moster. A medium power traveling-wave tube for 6,000-Mc radio relay. The Bell System Technical Journal, 35(6): 1285–1346, November 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-6-1285.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-6-
                                        1285.pdf.

Logan:1953:DME

[Log53] M. A. Logan. Dynamic measurements on electromagnetic devices. The Bell System Technical Journal, 32(6):1413–1467, November 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-6-1413.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-6-
                                        1413.pdf.

Logan:1954:ECO

[Log54] M. A. Logan. Estimation and control of the operate time of relays: Part II — design of optimum windings. The Bell System Technical Journal, 33(1):144–186, January 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-1-144.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-1-
144.pdf.

Louisell:1955:AST

[Lou55] W. H. Louisell. Analysis of the single tapered mode coupler. The Bell System Technical Journal, 34(4):853–870, July 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-4-853.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-4-
853.pdf.

Lozier:1951:ICN

[Loz51] J. C. Lozier. Instantaneous compandors on narrow band speech channels. The Bell System Technical Journal, 30(4):1214–1220, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1214.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1214.pdf.

Linvill:1956:DTT

[LS56] J. G. Linvill and L. G. Schimpf.

The design of tetrode transistor amplifiers. The Bell System Technical Journal, 35(4):813–840, July 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-4-813.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-4-
                                        813.pdf.

Lewis:1957:SDN

[LTLF57] H. A. Lewis, R. S. Tucker, G. H. Lovell, and J. M. Fraser. System design for the North Atlantic Link. The Bell System Technical Journal, 36(1):29–68, January 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol36/bstj36-
                                        1-29.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol36-1957/
                                        articles/bstj36-1-29.pdf.

Mason:1954:SSS

[MA54] W. P. Mason and O. L. Anderson. Stress systems in the solderless wrapped connection and their permanence. The Bell System Technical Journal, 33(5):1093– 1110, September 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-5-1093.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-5-
                                        1093.pdf.

MacColl:1951:REM

[Mac51] L. A. MacColl. On the reflection of electrons by metallic crystals. The Bell System Technical Journal, 30(4):888–906, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-888.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
888.pdf.

Mallinckrodt:1951:IC

[Mal51] C. O. Mallinckrodt. Instantaneous compandors. The Bell System Technical Journal, 30 (3):706–720, July 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-3-706.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-3-
706.pdf.

Mallina:1953:SWC

[Mal53] R. F. Mallina. Solderless wrapped connections: Part I — structure and tools. The Bell System Technical Journal, 32(3):525–555, May 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-3-525.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-3-
525.pdf.

Martin:1951:SFY

[Mar51] W. H. Martin. Seventy-five years of the telephone: An evolution in technology. The Bell System Technical Journal, 30(2):215–238, April 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-2-215.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-2-
                                        215.pdf.

Marcuse:1958:ATW

[Mar58] D. Marcuse. Attenuation of the TE01 wave within the curved helix waveguide. The Bell System Technical Journal, 37(6):1649– 1662, November 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-6-1649.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-6-
                                        1649.pdf.

Mason:1950:OPE

[Mas50] W. P. Mason. Optical properties and the electro-optic and photoelastic effects in crystals expressed in tensor form. The Bell System Technical Journal, 29(2):161–188, April 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-2-161.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-2-
                                        161.pdf.

Mason:1951:ZTC

[Mas51] W. P. Mason. Zero temperature coefficient quartz crystals for very high temperatures. The Bell System Technical Journal, 30 (2):366–380, April 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-2-366.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-2-
366.pdf.

Mason:1955:EDU

[Mas55] W. P. Mason. Effect of dislocations on ultrasonic wave attenuation in metals. The Bell System Technical Journal, 34(5):903–942, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-5-903.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-5-
903.pdf.

Maupin:1951:ISD

[Mau51] J. T. Maupin. Interaxial spacing and dielectric constant of pairs in multipaired cables. The Bell System Technical Journal, 30(3):652–667, July 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-3-652.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-3-
652.pdf.

McCluskey:1956:DGI

[McC56a] E. J. McCluskey, Jr. Detection of group invariance or total symmetry of a Boolean function. The Bell System Technical Journal, 35(6):1445–1453, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-6-1445.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-6-
1445.pdf.

McCluskey:1956:MBF

[McC56b] E. J. McCluskey, Jr. Minimization of Boolean functions. The Bell System Technical Journal, 35(6):1417–1444, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-6-1417.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-6-
1417.pdf.

McCall:1957:DPC

[McC57] D. W. McCall. The determination of pressure coefficients of capacitance for certain geometries. The Bell System Technical Journal, 36(2):485–495, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-2-485.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-2-
485.pdf.

McCluskey:1959:ECC

[McC59] E. J. McCluskey, Jr. Errorcorrecting codes — a linear programming approach. The Bell System Technical Journal, 38(6): 1485–1512, November 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1485.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1485.pdf.

McMillan:1952:IFRa

[McM52a] Brockway McMillan. Introduction to formal realizability theory — I. The Bell System Technical Journal, 31(2):217–279, March 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-2-217.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-2-
217.pdf.

McMillan:1952:IFRb

[McM52b] Brockway McMillan. Introduction to formal realizability theory — II. The Bell System Technical Journal, 31(3):541–600, May 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-3-541.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-3-
541.pdf.

McRae:1953:SWC

[McR53] J. W. McRae. Solderless wrapped connections — introduction. The Bell System Technical Journal, 32(3):523–524, May 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-3-523.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-3-
                                        523.pdf.

Mehring:1955:DVO

[ME55a] A. C. Mehring and E. L. Erwin. Dual voltage operation of relays and crossbar switches. The Bell System Technical Journal, 34(6):1225–1240, November 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-6-1225.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-6-
                                        1225.pdf.

Miller:1955:AJA

[ME55b] S. L. Miller and J. J. Ebers. Alloyed junction avalanche transistors. The Bell System Technical Journal, 34(5):883–902, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-5-883.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-5-
                                        883.pdf.

Mealy:1955:MSS

[Mea55] George H. Mealy. A method for synthesizing sequential circuits. The Bell System Technical Journal, 34(5):1045–1079, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-5-1045.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-5-
1045.pdf.

Menard:1952:NRM

[Men52] James Z. Menard. A new recording medium for transcribed message services. The Bell System Technical Journal, 31(3):530–540, May 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-3-530.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-3-
530.pdf.

Merrill:1951:TNI

[Mer51] J. L. Merrill, Jr. Theory of the negative impedance converter. The Bell System Technical Journal, 30 (1):88–109, January 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol30/bstj30-
1-88.pdf;   http://www.alcatel-
lucent.com/bstj/vol30-1951/
articles/bstj30-1-88.pdf.

Mottram:1957:TTC

[MHEG57] E. T. Mottram, R. J. Halsey, J. W. Emling, and R. G. Griffith. Transatlantic telephone cable system — planning and over-all performance. The Bell System Technical Journal, 36(1):7–27, January 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-1-7.pdf;   http://www.
alcatel-lucent.com/bstj/vol36-
1957/articles/bstj36-1-7.pdf.

Miller:1954:CWT

[Mil54a] S. E. Miller. Coupled wave theory and waveguide applications. The Bell System Technical Journal, 33(3):661–719, May 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-3-661.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-3-
661.pdf.

Miller:1954:WCM

[Mil54b] S. E. Miller. Waveguide as a communications medium. The Bell System Technical Journal, 33(6):1209–1265, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-6-1209.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-6-
1209.pdf.

Morris:1953:LCS

[MLD53] L. H. Morris, G. H. Lovell, and F. R. Dickinson. The L3 coaxial system: Amplifiers. The Bell System Technical Journal, 32(4):879–914, July 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-4-879.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-4-
879.pdf.

Mott:1951:RAT

[MM51] E. E. Mott and R. C. Miner.

The ring armature telephone receiver. The Bell System Technical Journal, 30(1):110–140, January 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-1-110.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-1-
110.pdf.

Mason:1952:MPP

[MM52] Warren P. Mason and H. J. Mc-Skimin. Mechanical properties of polymers at ultrasonic frequencies. The Bell System Technical Journal, 31(1):122–171, January 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-1-122.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-1-
122.pdf.

Mertz:1957:TAD

[MM57] P. Mertz and D. Mitchell. Transmission aspects of data transmission service using private line voice telephone channels. The Bell System Technical Journal, 36(6):1451–1486, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-6-1451.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-6-
                                        1451.pdf.

McNally:1957:ETT

[MMVH57] J. O. McNally, G. H. Metson, E. A. Veazie, and M. F. Holmes. Electron tubes for the transatlantic cable system. The Bell System Technical Journal, 36(1): 163–188, January 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-1-163.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-1-
                                        163.pdf.

Mason:1953:SWC

[MO53] W. P. Mason and T. F. Osmer. Solderless wrapped connections: Part II — necessary conditions for obtaining a permanent connection. The Bell System Technical Journal, 32(3):557–590, May 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/

bstj32-3-557.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-3-
557.pdf.

Montgomery:1952:ENS

[Mon52] H. C. Montgomery. Electrical noise in semiconductors. The Bell System Technical Journal, 31(5):950–975, September 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-950.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
950.pdf.

Morgan:1952:MTLb

[Mor52a] Samuel P. Morgan, Jr. Mathematical theory of laminated transmission lines — Part I. The Bell System Technical Journal, 31(5):883–949, September 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-883.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
883.pdf.

Morgan:1952:MTL

[Mor52b] Samuel P. Morgan, Jr. Mathematical theory of laminated transmission lines — Part II. The Bell System Technical Journal, 31(6):1121–1206, November 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-6-1121.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-6-
                                        1121.pdf.

Morton:1952:PST

[Mor52c] J. A. Morton. Present status of transistor development. The Bell System Technical Journal, 31(3):411–442, May 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-3-411.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-3-
                                        411.pdf.

Morgan:1957:TCC

[Mor57] Samuel P. Morgan. Theory of curved circular waveguide containing an inhomogeneous dielectric. The Bell System Technical Journal, 36(5):1209–1251, September 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-5-1209.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-5-
                                        1209.pdf.

Morin:1958:OTM

[Mor58] F. J. Morin. Oxides of the 3d transition metals. The Bell System Technical Journal, 37(4): 1047–1084, July 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol37/
bstj37-4-1047.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-4-
1047.pdf.

Meacham:1958:TRP

[MPW58] L. A. Meacham, J. R. Power, and F. West. Tone ringing and pushbutton calling. The Bell System Technical Journal, 37 (2):339–360, March 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-2-339.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-2-
339.pdf.

Morton:1950:DFB

[MR50] J. A. Morton and R. M. Ryder. Design factors of the Bell Telephone Laboratories 1553 triode. The Bell System Technical Journal, 29(4):496–530, October 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-4-496.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-4-
496.pdf.

Merrill:1954:NIT

[MRS54] J. L. Merrill, Jr., A. F. Rose, and J. O. Smethurst. Negative impedance telephone repeaters. The Bell System Technical Journal, 33(5):1055–1092, September 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-5-1055.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-5-
                                        1055.pdf.

Meszaros:1957:PFE

[MS57] G. W. Meszaros and H. H. Spencer. Power feed equipment for the North Atlantic Link. The Bell System Technical Journal, 36(1): 139–162, January 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-1-139.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-1-
                                        139.pdf.

Meitzler:1958:PDC

[MS58] A. H. Meitzler and H. L. Stadler. Piezoelectric and dielectric characteristics of single-crystal barium titanate plates. The Bell System Technical Journal, 37 (3):719–738, May 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-3-719.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-3-
                                        719.pdf.

Murphy:1958:CLI

[Mur58] R. B. Murphy. A criterion to limit inspection effort in continuous sampling plans. The Bell System Technical Journal, 37(1): 115–134, January 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-1-115.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-1-
115.pdf.

Malthaner:1952:EEC

[MV52] W. A. Malthaner and H. Earle Vaughan. An experimental electronically controlled automatic switching system. The Bell System Technical Journal, 31 (3):443–468, May 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-3-443.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-3-
443.pdf.

Mason:1952:NTM

[MW52] Warren P. Mason and Samuel D.

White. New techniques for measuring forces and wear in telephone switching apparatus. The Bell System Technical Journal, 31(3):469–503, May 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-3-469.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-3-
469.pdf.

Morgan:1956:HW

[MY56] S. P. Morgan and J. A. Young. Helix waveguide. The Bell System Technical Journal, 35(6):1347– 1384, November 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-6-1347.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-6-
                                        1347.pdf.

Myers:1952:CCT

[Mye52] Oscar Myers. Common control telephone switching systems. The Bell System Technical Journal, 31(6):1086–1120, November 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-6-1086.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-6-
                                        1086.pdf.

Narud:1959:ADT

[NA59] J. A. Narud and M. R. Aaron. Analysis and design of a transistor blocking oscillator including inherent nonlinearities. The Bell System Technical Journal, 38(3):785–852, May 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-3-785.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-3-
                                        785.pdf.

Nesenbergs:1959:LSS

[NM59] M. Nesenbergs and V. O. Mowery. Logic synthesis of some highspeed digital comparators. The Bell System Technical Journal, 38 (1):19–44, January 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol38/bstj38-
1-19.pdf;   http://www.alcatel-
lucent.com/bstj/vol38-1959/
articles/bstj38-1-19.pdf.

Nunn:1952:NNP

[Nun52] W. H. Nunn. Nationwide numbering plan. The Bell System Technical Journal, 31(5):851–859, September 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-851.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
851.pdf.

Ohl:1952:PIB

[Ohl52] Russell S. Ohl. Properties of ionic bombarded silicon. The Bell System Technical Journal, 31 (1):104–121, January 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-1-104.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-1-
104.pdf.

Oliver:1952:EC

[Oli52] B. M. Oliver. Efficient coding. The Bell System Technical Journal, 31(4):724–750, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-4-724.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
724.pdf.

Olmstead:1958:RDS

[Olm58] Paul S. Olmstead. Runs determined in a sample by an arbitrary cut. The Bell System Technical Journal, 37(1): 55–82, January 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol37/bstj37-
1-55.pdf;   http://www.alcatel-
lucent.com/bstj/vol37-1958/
articles/bstj37-1-55.pdf.

Owen:1959:MCB

[OW59] D. B. Owen and J. M. Wiesen. A method of computing bivariate normal probabilities: With an application to handling errors in testing and measuring. The Bell System Technical Journal, 38(2):553–572, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-553.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
553.pdf.

Purvis:1959:OPF

[PDH59] M. B. Purvis, G. V. Deverall, and D. R. Herriott. Optics and photography in the flying spot store. The Bell System Technical Journal, 38(2):403–424, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-403.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
403.pdf.

Peek:1951:ITR

[Pee51] R. L. Peek, Jr. Internal temperatures of relay windings. The Bell System Technical Journal, 30 (1):141–148, January 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-1-141.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-1-
141.pdf.

Peek:1954:AMM

[Pee54a] R. L. Peek, Jr. Analysis of measured magnetization and pull characteristics. The Bell System Technical Journal, 33(1): 79–108, January 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol33/bstj33-
1-79.pdf;   http://www.alcatel-
lucent.com/bstj/vol33-1954/
articles/bstj33-1-79.pdf.

Peek:1954:ECO

[Pee54b] R. L. Peek, Jr. Estimation and control of the operate time of relays: Part I — theory. The Bell System Technical Journal, 33 (1):109–143, January 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-1-109.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-1-
                                        109.pdf.

Peek:1954:PSR

[Pee54c] R. L. Peek, Jr. Principles of slow release relay design. The Bell System Technical Journal, 33 (1):187–217, January 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-1-187.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-1-
                                        187.pdf.

Pennoyer:1958:ANI

[Pen58] D. H. Pennoyer. Automatic number identification and its application to no. 1 crossbar, panel and step-by step offices. The Bell System Technical Journal, 37 (5):1295–1318, September 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-5-1295.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol37-1958/articles/bstj37-5-
1295.pdf.

Pferd:1954:GTD

[Pfe54] W. Pferd. A governor for telephone dials — principles of design. The Bell System Technical Journal, 33(6):1267–1307, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-6-1267.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-6-
1267.pdf.

Pierce:1950:TWTa

[Pie50a] J. R. Pierce. Traveling-wave tubes. The Bell System Technical Journal, 29(1):1–59, January 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol29-
1950/articles/bstj29-1-1.pdf.

Pierce:1950:TWTd

[Pie50b] J. R. Pierce. Traveling-wave tubes (fourth installment). The Bell System Technical Journal, 29(4): 608–671, October 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-4-608.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-4-
608.pdf.

Pierce:1950:TWTb

[Pie50c] J. R. Pierce. Traveling-wave tubes (second installment). The Bell System Technical Journal, 29 (2):189–250, April 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-2-189.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-2-
                                        189.pdf.

Pierce:1950:TWTc

[Pie50d] J. R. Pierce. Traveling-wave tubes (third installment). The Bell System Technical Journal, 29(3):390–460, July 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-3-390.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-3-
                                        390.pdf.

Pierce:1951:GSE

[Pie51a] J. R. Pierce. A gun for starting electrons straight in a magnetic field. The Bell System Technical Journal, 30(4):825–829, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-4-825.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-4-
                                        825.pdf.

Pierce:1951:WES

[Pie51b] J. R. Pierce. Waves in electron streams and circuits. The Bell System Technical Journal, 30(3):626–651, July 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-3-626.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-3-
626.pdf.

Pierce:1954:WPM

[Pie54] J. R. Pierce. The wave picture of microwave tubes. The Bell System Technical Journal, 33(6):1343–1372, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-6-1343.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-6-
1343.pdf.

Pilliod:1952:FPT

[Pil52] J. J. Pilliod. Fundamental plans for toll telephone plant. The Bell System Technical Journal, 31(5):832–850, September 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-832.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
832.pdf.

Pierce:1957:RRI

[PK57] J. R. Pierce and J. E. Karlin. Reading rates and the information rate of a human channel. The Bell System Technical Journal, 36(2):497–516, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-2-497.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-2-
497.pdf.

Prim:1951:SRC

[Pri51] R. C. Prim, III. Some results concerning the partial differential equations describing the flow of holes and electrons in semiconductors. The Bell System Technical Journal, 30(4):1174–1213, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1174.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1174.pdf.

Prim:1953:DFD

[Pri53] R. C. Prim. DC field distribution in a ‘swept intrinsic’ semiconductor configuration. The Bell System Technical Journal, 32(3):665–694, May 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-3-665.pdf;   http://
www.alcatel-lucent.com/bstj/

vol32-1953/articles/bstj32-3-
665.pdf.

Prince:1956:DPJ

[Pri56] M. B. Prince. Diffused p–n junction silicon rectifiers. The Bell System Technical Journal, 35(3):661–684, May 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-3-661.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-3-
661.pdf.

Prim:1957:SCN

[Pri57] R. C. Prim. Shortest connection networks and some generalizations. The Bell System Technical Journal, 36(6):1389–1401, November 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-6-1389.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-6-
1389.pdf.

Peek:1954:MDR

[PW54] R. L. Peek, Jr. and H. N. Wagar. Magnetic design of relays. The Bell System Technical Journal, 33(1):23–78, January 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol33/bstj33-
1-23.pdf;   http://www.alcatel-
lucent.com/bstj/vol33-1954/
articles/bstj33-1-23.pdf.

Pierce:1956:GWD

[PW56] J. R. Pierce and L. R. Walker. Growing waves due to transverse velocities. The Bell System Technical Journal, 35(1): 109–125, January 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-1-109.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-1-
                                        109.pdf.

Quinlan:1954:ACW

[Qui54] A. L. Quinlan. Automatic contact welding in wire spring relay manufacture. The Bell System Technical Journal, 33(4):897–923, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-4-897.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-4-
                                        897.pdf.

Rafuse:1954:NMR

[Raf54] I. S. Rafuse. A new multicontact relay for telephone switching systems. The Bell System Technical Journal, 33(5):1111–1132, September 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-5-1111.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-5-
                                        1111.pdf.

Raisbeck:1958:ACL

[Rai58] Gordon Raisbeck. Attenuation in continuously loaded coaxial cables. The Bell System Technical Journal, 37(2):361–374, March 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-2-361.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-2-
361.pdf.

Raisbeck:1959:NLT

[Rai59] Gordon Raisbeck. Nonuniformities in laminated transmission lines. The Bell System Technical Journal, 38(2):477–516, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-2-477.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-2-
477.pdf.

Rausch:1959:AVC

[Rau59] R. G. Rausch. The analysis of valve-controlled hydraulic servomechanisms. The Bell System Technical Journal, 38(6):1513– 1549, November 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1513.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1513.pdf.

Read:1956:TSI

[Rea56] W. T. Read, Jr. Theory of the swept intrinsic structure. The Bell System Technical Journal, 35(6):1239–1284, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-6-1239.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-6-
                                        1239.pdf.

Read:1958:PHF

[Rea58] W. T. Read, Jr. A proposed high-frequency, negative-resistance diode. The Bell System Technical Journal, 37(2):401–446, March 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-2-401.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-2-
                                        401.pdf.

Reed:1953:CRR

[Ree53] E. D. Reed. A coupled resonator reflex klystron. The Bell System Technical Journal, 32(3):715–766, May 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-3-715.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-3-
                                        715.pdf.

Reed:1955:TLV

[Ree55] E. D. Reed. A tunable, lowvoltage reflex klystron for operation in the 50 to 60-kmc band. The Bell System Technical Journal, 34(3):563–599, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-3-563.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-3-
563.pdf.

Reiss:1956:CIA

[RFM56] Howard Reiss, C. S. Fuller, and F. J. Morin. Chemical interactions among defects in germanium and silicon. The Bell System Technical Journal, 35(3):535–636, May 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-3-535.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-3-
535.pdf.

Rieke:1953:LCS

[RG53] John W. Rieke and R. S. Graham.

The L3 coaxial system: Television terminals. The Bell System Technical Journal, 32(4):915–942, July 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-4-915.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-4-
915.pdf.

Rice:1950:CPN

[Ric50a] S. O. Rice. Communication in the presence of noise — probability of error for two encoding schemes. The Bell System Technical Journal, 29(1):60–93, January 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol29/bstj29-
                                        1-60.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol29-1950/
                                        articles/bstj29-1-60.pdf.

Richardson:1950:LTF

[Ric50b] J. M. Richardson. The linear theory of fluctuations arising from diffusional mechanisms — an attempt at a theory of contact noise. The Bell System Technical Journal, 29(1):117–141, January 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-1-117.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-1-
                                        117.pdf.

Rice:1954:DPR

[Ric54] S. O. Rice. Diffraction of plane radio waves by a parabolic cylinder. The Bell System Technical Journal, 33(2):417–504, March 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-2-417.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-2-
                                        417.pdf.

Rice:1957:DPN

[Ric57] S. O. Rice. Distortion produced in a noise modulated FM signal by nonlinear attenuation and phase shift. The Bell System Technical Journal, 36(4):879–889, July 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-4-879.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-4-
879.pdf.

Rice:1958:DDF

[Ric58] S. O. Rice. Distribution of the duration of fades in radio transmission: Gaussian noise model. The Bell System Technical Journal, 37(3):581–635, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-581.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-3-
581.pdf.

Rice:1959:RTB

[Ric59] L. P. Rice. Radio transmission into buildings at 35 and 150 mc. The Bell System Technical Journal, 38(1):197–210, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-1-197.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-1-
197.pdf.

Rigrod:1957:NSEb

[Rig57a] W. W. Rigrod. Noise spectrum of electron beam in longitudal magnetic field: Part II — the UHF noise spectrum. The Bell System Technical Journal, 36(4):855–878, July 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-4-855.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-4-
                                        855.pdf.

Rigrod:1957:NSEa

[Rig57b] W. W. Rigrod. Noise spectrum of electron beam in longitudinal magnetic field: Part I — the growing noise phenomenon. The Bell System Technical Journal, 36(4):831–853, July 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-4-831.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-4-
                                        831.pdf.

Rigrod:1959:SCWb

[Rig59] W. W. Rigrod. Space-charge wave harmonics and noise propagation in rotating electron beams. The Bell System Technical Journal, 38(1):119–139, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-1-119.pdf;   http://

www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-1-
119.pdf.

Riordan:1951:TTT

[Rio51] John Riordan. Telephone traffic time average. The Bell System Technical Journal, 30(4): 1129–1144, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1129.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1129.pdf.

Riordan:1953:ADC

[Rio53a] John Riordan. Addendum to Delay curves for calls served at random. The Bell System Technical Journal, 32(5):1266, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-5-1266.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-5-
1266.pdf.

Riordan:1953:DCC

[Rio53b] John Riordan. Delay curves for calls served at random. The Bell System Technical Journal, 32 (1):100–119, January 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-1-100.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol32-1953/articles/bstj32-1-
                                        100.pdf.

Rigrod:1954:WPA

[RL54] W. W. Rigrod and J. A. Lewis. Wave propagation along a magneticallyfocused cylindrical electron beam. The Bell System Technical Journal, 33(2):399–416, March 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-2-399.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-2-
                                        399.pdf.

Rounds:1955:ECL

[RL55] P. W. Rounds and Miss G. L. Lakin. Equalization of cables for local television transmission. The Bell System Technical Journal, 34(4):713–738, July 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-4-713.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-4-
                                        713.pdf.

Raisbeck:1958:TCT

[RM58] G. Raisbeck and J. M. Manley. Transmission characteristics of a three-conductor coaxial transmission line with transpositions. The Bell System Technical Journal, 37(4):835–876, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol37/
bstj37-4-835.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-4-
835.pdf.

Robb:1956:ATT

[Rob56] D. T. Robb. Automatic testing in telephone manufacture. The Bell System Technical Journal, 35(5):1129–1154, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-5-1129.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-5-
1129.pdf.

Roberts:1958:PCC

[Rob58] S. W. Roberts. Properties of control chart zone tests. The Bell System Technical Journal, 37 (1):83–114, January 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol37/bstj37-
1-83.pdf;   http://www.alcatel-
lucent.com/bstj/vol37-1958/
articles/bstj37-1-83.pdf.

Rose:1958:RMH

[Ros58] C. F. P. Rose. Research models of helix waveguides. The Bell System Technical Journal, 37(3):679–688, May 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-679.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-3-
                                        679.pdf.

Rowen:1953:FMA

[Row53] J. H. Rowen. Ferrites in microwave applications. The Bell System Technical Journal, 32(6): 1333–1369, November 1953. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-6-1333.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-6-
                                        1333.pdf.

Rowe:1958:TLC

[Row58] H. E. Rowe. Timing in a long chain of regenerative binary repeaters. The Bell System Technical Journal, 37(6):1543–1598, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-6-1543.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-6-
                                        1543.pdf.

Rigrod:1959:SCWa

[RP59] W. W. Rigrod and J. R. Pierce. Space-charge wave excitation in solid-cylindrical Brillouin beams. The Bell System Technical Journal, 38(1):99–118, January 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol38/bstj38-
                                        1-99.pdf;   http://www.alcatel-

lucent.com/bstj/vol38-1959/
articles/bstj38-1-99.pdf.

Roetken:1951:TMR

[RSF51] A. A. Roetken, K. D. Smith, and R. W. Friis. The TD-2 microwave radio relay system. The Bell System Technical Journal, 30(4):1041–1077, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1041.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1041.pdf.

Ruthroff:1958:AMS

[Rut58] C. L. Ruthroff. Amplitude modulation suppression in FM systems. The Bell System Technical Journal, 37(4):1023–1046, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-4-1023.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-4-
1023.pdf.

Seidel:1958:FSC

[SB58] H. Seidel and H. Boyet. Frequency shifts in cavities with longitudinally magnetized small ferrite discs. The Bell System Technical Journal, 37(3):637–655, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-637.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-3-
                                        637.pdf.

Smith:1959:FPS

[SB59] J. L. Smith and W. S. Boyle. Fundamental processes of the short arc: With applications to contact erosion and percussion welding. The Bell System Technical Journal, 38(2):537–552, March 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-2-537.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-2-
                                        537.pdf.

Schneckloth:1951:SBC

[Sch51] H. H. Schneckloth. Some basic concepts of translators and identifiers used in telephone switching systems. The Bell System Technical Journal, 30(3):588–625, July 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol30/
                                        bstj30-3-588.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol30-1951/articles/bstj30-3-
                                        588.pdf.

Schelkunoff:1952:GTE

[Sch52] S. A. Schelkunoff. Generalized telegraphist’s equations for waveguides. The Bell System Technical Journal, 31(4):784–801, July 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol31/
bstj31-4-784.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-4-
784.pdf.

Schelkunoff:1955:CME

[Sch55] S. A. Schelkunoff. Conversion of Maxwell’s equations into generalized telegraphist’s equations. The Bell System Technical Journal, 34(5):995–1043, September 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-5-995.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-5-
995.pdf.

Schaefer:1957:EOH

[Sch57] J. W. Schaefer. An electrically operated hydraulic control valve. The Bell System Technical Journal, 36(3):711–736, May 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-3-711.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-3-
711.pdf.

Schulz-DuBois:1959:PSS

[SD59] E. O. Schulz-DuBois. Paramagnetic spectra of substituted sapphires–Part I: Ruby. The Bell System Technical Journal, 38(1): 271–290, January 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-1-271.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-1-
                                        271.pdf.

Schulz-DuBois:1959:UAM

[SDSD59] E. O. Schulz-DuBois, H. E. D. Scovil, and R. W. DeGrasse. Use of active material in threelevel solid state masers. The Bell System Technical Journal, 38 (2):335–352, March 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol38/
                                        bstj38-2-335.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol38-1959/articles/bstj38-2-
                                        335.pdf.

Seidel:1957:CWM

[Sei57] H. Seidel. The character of waveguide modes in gyromagnetic media. The Bell System Technical Journal, 36(2):409–426, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-2-409.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-2-
                                        409.pdf.

Semon:1958:SSP

[Sem58] Warren Semon. Synthesis of seriesparallel network switching functions. The Bell System Technical Journal, 37(4):877–898, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-4-877.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-4-
877.pdf.

Seidel:1959:GMW

[SF59] H. Seidel and R. C. Fletcher.

Gyromagnetic modes in waveguide partially loaded with ferrite. The Bell System Technical Journal, 38(6):1427–1456, November 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1427.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1427.pdf.

Sobel:1959:GTE

[SG59] Milton Sobel and Phyllis A. Groll.

Group testing to eliminate efficiently all defectives in a binomial sample. The Bell System Technical Journal, 38(5):1179–1252, September 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-5-1179.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-5-
1179.pdf.

Sobel:1957:SBO

[SH57] Milton Sobel and Marilyn J.

Huyett. Selecting the best one of several binomial populations. The Bell System Technical Journal, 36(2):537–576, March 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-2-537.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-2-
                                        537.pdf.

Sobel:1958:NDR

[SH58] Milton Sobel and Marilyn J. Huyett. Nonparametric definition of the representativeness of a sample — with tables. The Bell System Technical Journal, 37 (1):135–161, January 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-1-135.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-1-
                                        135.pdf.

Shannon:1950:MRT

[Sha50] Claude E. Shannon. Memory requirements in a telephone exchange. The Bell System Technical Journal, 29(3):343–349, July 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-3-343.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-3-
                                        343.pdf.

Shannon:1951:PEP

[Sha51a] C. E. Shannon. Prediction and entropy of printed English. The Bell System Technical Journal, 30 (1):50–64, January 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol30/bstj30-
1-50.pdf;   http://www.alcatel-
lucent.com/bstj/vol30-1951/
articles/bstj30-1-50.pdf.

Shaw:1951:EILa

[Sha51b] Thomas Shaw. The evolution of inductive loading for Bell System telephone facilities. The Bell System Technical Journal, 30(1): 149–204, January 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-1-149.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-1-
149.pdf.

Shaw:1951:EILc

[Sha51c] Thomas Shaw. The evolution of inductive loading for Bell System telephone facilities. The Bell System Technical Journal, 30(3):721–764, July 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-3-721.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-3-
721.pdf.

Shaw:1951:EILd

[Sha51d] Thomas Shaw. The evolution of inductive loading for Bell System telephone facilities. The Bell System Technical Journal, 30(4): 1221–1243, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1221.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1221.pdf.

Shaw:1951:EILb

[Sha51e] Thomas Shaw. The evolution of inductive loading for Bell System telephone facilities Part III: Loading for exchange area cables. The Bell System Technical Journal, 30(2):447–472, April 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-2-447.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-2-
447.pdf.

Sharpless:1956:WTM

[Sha56] W. M. Sharpless. Wafer-type millimeter wave rectifiers. The Bell System Technical Journal, 35(6):1385–1402, November 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-6-1385.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-6-
1385.pdf.

Shannon:1959:PEO

[Sha59a] Claude E. Shannon. Probability of error for optimal codes in a Gaussian channel. The Bell System Technical Journal, 38(3):611–656, May 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-3-611.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-3-
611.pdf.

Sharpless:1959:HFG

[Sha59b] W. M. Sharpless. High-frequency gallium arsenide point-contact rectifiers. The Bell System Technical Journal, 38(1):259–269, January 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-1-259.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-1-
259.pdf.

Shewhart:1958:NOS

[She58] W. A. Shewhart. Nature and origin of standards of quality. The Bell System Technical Journal, 37(1):1–22, January 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol37-
1958/articles/bstj37-1-1.pdf.

Shipley:1952:ATS

[Shi52] F. F. Shipley. Automatic toll switching systems. The Bell System Technical Journal, 31(5): 860–882, September 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-5-860.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-5-
860.pdf.

Shockley:1951:HEG

[Sho51] W. Shockley. Hot electrons in germanium and Ohm’s Law. The Bell System Technical Journal, 30 (4):990–1034, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-990.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
990.pdf.

Shockley:1954:NRA

[Sho54] W. Shockley. Negative resistance arising from transit time in semiconductor diodes. The Bell System Technical Journal, 33(4):799–826, July 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-799.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
799.pdf.

Slepian:1956:CBS

[Sle56] David Slepian. A class of binary signaling alphabets. The Bell System Technical Journal, 35 (1):203–234, January 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-1-203.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-1-
203.pdf.

Slepian:1958:FRN

[Sle58] D. Slepian. Fluctuations of random noise power. The Bell System Technical Journal, 37(1): 163–184, January 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-1-163.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-1-
163.pdf.

Smith:1957:ICQ

[Smi57] Bernard Smith. Instantaneous companding of quantized signals. The Bell System Technical Journal, 36(3):653–709, May 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-3-653.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-3-
653.pdf.

Smits:1958:MSR

[Smi58] F. M. Smits. Measurement of sheet resistivities with the fourpoint probe. The Bell System Technical Journal, 37(3):711–718, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-711.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-3-
711.pdf.

Snoke:1957:ROM

[Sno57] Lloyd R. Snoke. Resistance of organic materials and cable structures to marine biological attack. The Bell System Technical Journal, 36(5):1095–1127, September 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-5-1095.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-5-
1095.pdf.

Sobel:1956:STR

[Sob56] Milton Sobel. Statistical techniques for reducing the experiment time in reliability studies. The Bell System Technical Journal, 35(1):179–202, January 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-1-179.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-1-
179.pdf.

Southworth:1950:PAW

[Sou50] George C. Southworth. Principles and applications of waveguide transmission. The Bell System Technical Journal, 29 (3):295–342, July 1950. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol29/
bstj29-3-295.pdf;   http://
www.alcatel-lucent.com/bstj/
vol29-1950/articles/bstj29-3-
295.pdf.

Staehler:1952:ABA

[Sta52] Robert E. Staehler. An application of Boolean algebra to switching circuit design. The Bell System Technical Journal, 31 (2):280–305, March 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-2-280.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-2-
280.pdf.

Stone:1953:FCI

[Sto53] H. A. Stone, Jr. Ferrite core inductors. The Bell System Technical Journal, 32(2):265–291, March 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-2-265.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-2-
265.pdf.

Stubner:1953:AEE

[Stu53] F. W. Stubner. Acceleration effects on electron tubes. The Bell System Technical Journal, 32 (5):1203–1229, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-5-1203.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-5-
                                        1203.pdf.

Suhl:1953:TME

[Suh53] Harry Suhl. Theory of magnetic effects on the noise in a germanium filament. The Bell System Technical Journal, 32(3):647–664, May 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-3-647.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-3-
                                        647.pdf.

Sumner:1952:RAR

[Sum52] Eric Eden Sumner. Relay armature rebound analysis. The Bell System Technical Journal, 31 (1):172–200, January 1952. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-1-172.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-1-
                                        172.pdf.

Sumner:1954:SFP

[Sum54] Eric Eden Sumner. Some fundamental problems in percussive welding. The Bell System Technical Journal, 33(4):885–895, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-885.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
885.pdf.

Sunde:1954:TFPa

[Sun54a] E. D. Sunde. Theoretical fundamentals of pulse transmission — I. The Bell System Technical Journal, 33(3):721–788, May 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-3-721.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-3-
721.pdf.

Sunde:1954:TFPb

[Sun54b] E. D. Sunde. Theoretical fundamentals of pulse transmission — II. The Bell System Technical Journal, 33(4):987–1010, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-987.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
987.pdf.

Sunde:1957:STR

[Sun57] E. D. Sunde. Self-timing regenerative repeaters. The Bell System Technical Journal, 36 (4):891–937, July 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-4-891.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-4-
891.pdf.

Sunde:1959:IBP

[Sun59] E. D. Sunde. Ideal binary pulse transmission by AM and FM. The Bell System Technical Journal, 38(6):1357–1426, November 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1357.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1357.pdf.

Suhl:1954:TGWa

[SW54a] H. Suhl and L. R. Walker. Topics in guided-wave propagation through gyromagnetic media: Part I — the completely filled cylindrical guide. The Bell System Technical Journal, 33(3):579–659, May 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-3-579.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-3-
579.pdf.

Suhl:1954:TGWb

[SW54b] H. Suhl and L. R. Walker. Topics in guided wave propagation through gyromagnetic media: Part II — transverse magnetization and the non-reciprocal helix. The Bell System Technical Journal, 33(4):939–986, July 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-4-939.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-4-
939.pdf.

Suhl:1954:TGWc

[SW54c] H. Suhl and L. R. Walker. Topics in guided wave propagation through gyromagnetic media: Part III — perturbation theory and miscellaneous results. The Bell System Technical Journal, 33(5): 1133–1194, September 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-5-1133.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-5-
1133.pdf.

Seckler:1958:FDS

[SY58] Howard N. Seckler and John J.

Yostpille. Functional design of a stored-program electronic switching system. The Bell System Technical Journal, 37(6):1327– 1382, November 1958. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-6-1327.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-6-
                                        1327.pdf.

Townsend:1953:CCT

[TD53] M. A. Townsend and W. A. Depp. Cold cathode tubes for transmission of audio frequency signals. The Bell System Technical Journal, 32(6):1371–1391, November 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-6-1371.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-6-
                                        1371.pdf.

Tendick:1956:TPR

[Ten56] F. H. Tendick, Jr. Transistor pulse regenerative amplifiers. The Bell System Technical Journal, 35(5):1085–1114, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-1085.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1085.pdf.

Thomas:1956:TPS

[Tho56] D. E. Thomas. Tables of phase of a semi-infinite unit attenuation slope. The Bell System Technical Journal, 35(3):747–749, May 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-3-747.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-3-
747.pdf.

Thomas:1959:SDC

[Tho59] D. E. Thomas. Some design considerations for high-frequency transistor amplifiers. The Bell System Technical Journal, 38(6): 1551–1580, November 1959. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1551.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1551.pdf.

Tien:1956:LST

[Tie56] Ping King Tien. A large signal theory of traveling wave amplifiers: Including the effects of space charge and finite coupling between the beam and the circuit. The Bell System Technical Journal, 35(2):349–374, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-2-349.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-2-
349.pdf.

Thomas:1957:PFS

[TK57] J. F. P. Thomas and R. Kelly.

Power-feed system for the Newfoundland– Nova Scotia link. The Bell System Technical Journal, 36(1): 277–292, January 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-1-277.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-1-
                                        277.pdf.

Townsend:1957:CCG

[Tow57] M. A. Townsend. Cold cathode gas tubes for telephone switching systems. The Bell System Technical Journal, 36(3):755–768, May 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-3-755.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-3-
                                        755.pdf.

Truitt:1954:TET

[Tru54] C. J. Truitt. Traffic engineering techniques for determining trunk requirements in alternate routing trunk networks. The Bell System Technical Journal, 33 (2):277–302, March 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-2-277.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-2-
                                        277.pdf.

Tanenbaum:1956:DEB

[TT56] M. Tanenbaum and D. E. Thomas.

Diffused emitter and base silicon transistors. The Bell System Technical Journal, 35(1):1–22, January 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol35-
1956/articles/bstj35-1-1.pdf.

Uhlir:1955:PIS

[Uhl55] Arthur Uhlir, Jr. The potentials of infinite systems of sources and numerical solutions of problems in semi-conductor engineering. The Bell System Technical Journal, 34(1):105–128, January 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol34/
bstj34-1-105.pdf;   http://
www.alcatel-lucent.com/bstj/
vol34-1955/articles/bstj34-1-
105.pdf.

Uhlir:1956:ESG

[Uhl56] A. Uhlir, Jr. Electrolytic shaping of germanium and silicon. The Bell System Technical Journal, 35(2):333–347, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-2-333.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-2-
333.pdf.

Uhlir:1958:SNP

[Uhl58] A. Uhlir, Jr. Shot noise in p–n junction frequency converters. The Bell System Technical Journal, 37(4):951–988, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-4-951.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-4-
                                        951.pdf.

Ulrich:1957:NBE

[Ulr57] Werner Ulrich. Non-binary error correction codes. The Bell System Technical Journal, 36(6): 1341–1388, November 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-6-1341.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-6-
                                        1341.pdf.

Unger:1957:CEWa

[Ung57a] H. G. Unger. Circular electric wave transmission in a dielectriccoated waveguide. The Bell System Technical Journal, 36(5):1253– 1278, September 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-5-1253.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-5-
                                        1253.pdf.

Unger:1957:CEWb

[Ung57b] H. G. Unger. Circular electric wave transmission through serpentine bends. The Bell System Technical Journal, 36(5):1279–1291, September 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-5-1279.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-5-
1279.pdf.

Unger:1957:NMB

[Ung57c] H. G. Unger. Normal mode bends for circular electric waves. The Bell System Technical Journal, 36(5):1292–1307, September 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-5-1292.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-5-
1292.pdf.

Unger:1958:CWT

[Ung58a] Hans-Georg Unger. Circular waveguide taper of improved design. The Bell System Technical Journal, 37(4):899–912, July 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-4-899.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-4-
899.pdf.

Unger:1958:HWT

[Ung58b] Hans-Georg Unger. Helix waveguide theory and application. The Bell System Technical Journal, 37(6):1599–1647, November 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-6-1599.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol37-1958/articles/bstj37-6-
                                        1599.pdf.

Vaage:1953:TPL

[Vaa53] E. F. Vaage. Transmission properties of laminated clogston type conductors. The Bell System Technical Journal, 32(3):695–713, May 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol32/
                                        bstj32-3-695.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol32-1953/articles/bstj32-3-
                                        695.pdf.

VanRoosbroeck:1950:TFE

[Van50] W. Van Roosbroeck. Theory of the flow of electrons and holes in germanium and other semiconductors. The Bell System Technical Journal, 29(4):560–607, October 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol29/
                                        bstj29-4-560.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol29-1950/articles/bstj29-4-
                                        560.pdf.

VanHorn:1953:SWC

[Van53] R. H. Van Horn. Solderless wrapped connections: Part III — evaluation and performance tests. The Bell System Technical Journal, 32(3):591–610, May 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-3-591.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-3-
591.pdf.

Vaughan:1959:RMT

[Vau59] H. E. Vaughan. Research model for time-separation integrated communication. The Bell System Technical Journal, 38(4):909–932, July 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-4-909.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-4-
909.pdf.

Veloric:1957:HVC

[VP57] H. S. Veloric and M. B. Prince.

High-voltage conductivity-modulated silicon rectifier. The Bell System Technical Journal, 36(4): 975–1004, July 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-4-975.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-4-
975.pdf.

VonAulock:1957:MDM

[VR57] Wilhelm Von Aulock and John H.

Rowen. Measurement of dielectric and magnetic properties of ferromagnetic materials at microwave frequencies. The Bell System Technical Journal, 36(2): 427–448, March 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-2-427.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-2-
                                        427.pdf.

Wagar:1954:ETR

[Wag54a] H. N. Wagar. Economics of telephone relay applications. The Bell System Technical Journal, 33 (1):218–256, January 1954. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-1-218.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol33-1954/articles/bstj33-1-
                                        218.pdf.

Wagar:1954:RME

[Wag54b] H. N. Wagar. Relay measuring equipment. The Bell System Technical Journal, 33(1):3–22, January 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol33/
                                        bstj33-1-3.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol33-
                                        1954/articles/bstj33-1-3.pdf.

Wallace:1951:RMR

[Wal51] R. L. Wallace, Jr. The reproduction of magnetically recorded signals. The Bell System Technical Journal, 30(4):1145–1173, October 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-1145.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
1145.pdf.

Wannier:1953:MGI

[Wan53] Gregory H. Wannier. Motion of gaseous ions in strong electric fields. The Bell System Technical Journal, 32(1):170–254, January 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-1-170.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-1-
170.pdf.

Warters:1958:EMF

[War58] W. D. Warters. The effects of mode filters on the transmission characteristics of circular electric waves in a circular waveguide. The Bell System Technical Journal, 37(3):657–677, May 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol37/
bstj37-3-657.pdf;   http://
www.alcatel-lucent.com/bstj/
vol37-1958/articles/bstj37-3-
657.pdf.

Weissmann:1957:SRR

[Wei57] G. F. Weissmann. Strength requirements for round conduit. The Bell System Technical Journal, 36(3):737–754, May 1957. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol36/
                                        bstj36-3-737.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol36-1957/articles/bstj36-3-
                                        737.pdf.

Welber:1955:PST

[WEP55] I. Welber, H. W. Evans, and G. A. Pullis. Protection of service in the TD-2 radio relay system by automatic channel switching. The Bell System Technical Journal, 34(3):473–510, May 1955. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol34/
                                        bstj34-3-473.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol34-1955/articles/bstj34-3-
                                        473.pdf.

Wilhelm:1952:IBM

[Wil52] H. T. Wilhelm. Impedance bridges for the megacycle range. The Bell System Technical Journal, 31(5):999–1012, September 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol31/
                                        bstj31-5-999.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-5-
                                        999.pdf.

Wilkinson:1953:WCD

[Wil53] Roger I. Wilkinson. Working curves for delayed exponential calls served in random order. The Bell System Technical Journal, 32(2):360–383, March 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-2-360.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-2-
360.pdf.

Wilkinson:1956:TTT

[Wil56] Roger I. Wilkinson. Theories for toll traffic engineering in the U.S.A. The Bell System Technical Journal, 35(2):421–514, March 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol35/
bstj35-2-421.pdf;   http://
www.alcatel-lucent.com/bstj/
vol35-1956/articles/bstj35-2-
421.pdf.

Windeler:1953:PIT

[Win53] A. S. Windeler. Polyethylene insulated telephone cable. The Bell System Technical Journal, 32 (5):1245–1256, September 1953. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol32/
bstj32-5-1245.pdf;   http://
www.alcatel-lucent.com/bstj/
vol32-1953/articles/bstj32-5-
1245.pdf.

Weissmann:1959:ETP

[WM59] G. F. Weissmann and Duncan M. Mitchel. Effects of tamping and pavement breaking on round conduit. The Bell System Technical Journal, 38(6):1581–1594, November 1959. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol38/
bstj38-6-1581.pdf;   http://
www.alcatel-lucent.com/bstj/
vol38-1959/articles/bstj38-6-
1581.pdf.

Weaver:1954:BSF

[WN54] A. Weaver and N. A. Newell. Inband single-frequency signaling. The Bell System Technical Journal, 33(6):1309–1330, November 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-6-1309.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-6-
1309.pdf.

Wojciechowski:1954:CIT

[Woj54] B. M. Wojciechowski. Continuous incremental thickness measurements of non-conductive cable sheath. The Bell System Technical Journal, 33(2):353–368, March 1954. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol33/
bstj33-2-353.pdf;   http://
www.alcatel-lucent.com/bstj/
vol33-1954/articles/bstj33-2-
353.pdf.

Wood:1951:DFC

[Woo51] Elizabeth A. Wood. Detwinning ferroelectric crystals. The Bell System Technical Journal, 30(4): 945–955, October 1951. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-4-945.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-4-
945.pdf.

Wallace:1951:SCP

[WP51] R. L. Wallace, Jr. and W. J.

Pietenpol. Some circuit properties and applications of n–p–n transistors. The Bell System Technical Journal, 30(3):530–563, July 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-3-530.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-3-
530.pdf.

Wallace:1951:DGT

[WR51] R. L. Wallace, Jr. and G. Raisbeck.

Duality as a guide in transistor circuit design. The Bell System Technical Journal, 30(2):381–417, April 1951. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol30/
bstj30-2-381.pdf;   http://
www.alcatel-lucent.com/bstj/
vol30-1951/articles/bstj30-2-
381.pdf.

Wrathall:1956:TBP

[Wra56] L. R. Wrathall. Transistorized binary pulse regenerator. The Bell System Technical Journal, 35 (5):1059–1084, September 1956. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-5-1059.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-5-
                                        1059.pdf.

Weisbaum:1956:FDI

[WS56] S. Weisbaum and H. Seidel. The field displacement isolator. The Bell System Technical Journal, 35(4):877–898, July 1956. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol35/
                                        bstj35-4-877.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol35-1956/articles/bstj35-4-
                                        877.pdf.

Yokelson:1958:SCD

[YCU58] B. J. Yokelson, W. B. Cagle, and M. D. Underwood. Semiconductor circuit design philosophy for the central control of an electronic switching system. The Bell System Technical Journal, 37(5):1125–1160, September 1958. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol37/
                                        bstj37-5-1125.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol37-1958/articles/bstj37-5-
1125.pdf.

Young:1952:CMR

[You52] W. Rae Young, Jr. Comparison of mobile radio transmission at 150, 450, 900, and 3700 Mc. The Bell System Technical Journal, 31(6):1068–1085, November 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-6-1068.pdf;   http://
www.alcatel-lucent.com/bstj/
vol31-1952/articles/bstj31-6-
1068.pdf.

Zajac:1957:DKL

[Zaj57] E. E. Zajac. Dynamics and kinematics of the laying and recovery of submarine cable. The Bell System Technical Journal, 36 (5):1129–1207, September 1957. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol36/
bstj36-5-1129.pdf;   http://
www.alcatel-lucent.com/bstj/
vol36-1957/articles/bstj36-5-
1129.pdf.

Zinn:1952:NRT

[Zin52] M. K. Zinn. Network representation of transcendental impedance functions. The Bell System Technical Journal, 31(2):378–404, March 1952. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol31/
bstj31-2-378.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol31-1952/articles/bstj31-2-
                                        378.pdf.